A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1960–1964

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.15

Title word cross-reference

#1023 [740].

(r^{−4}) [117]. 1/2 [536]. 2 [606]. 3 + 2 [434]. 4 [618]. 8 [618]. [□^{m} − (−μ^{2})^{m}]^{l} [321]. E [451]. F_{1} [562]. F_{D} [562]. H [87, 220]. J_{0}(x) [651]. k [156]. l [451]. m [603]. N [382, 377, 622, 728, 433, 689, 584, 503, 211]. N/D [685]. p [153]. R [479]. R_{4} [110]. R_{8} [435]. S [205, 79, 451, 566, 173, 732, 80, 275, 490, 421]. SU(3) [673, 738, 547]. SU_{3} [726, 632, 578]. SU_{n} [725]. T [447, 392].

-Body [382, 584, 503]. -Commutative [156]. -Dimensional [606, 618, 433, 689]. -Identical [211]. -Matrix [490, 392, 479, 275, 421]. -Particle [622, 728].

2-Semi-Isomorphic [129]. 26 [740].

4 [740].

Absolute [100, 266]. Absorption [159, 160, 177]. Accuracy [539]. Acnodes [134, 430]. Action [174, 625]. Action-at-a-Distance [174]. Active [129]. Adapted [51, 141]. Addendum [733, 220, 470]. Addition [543]. Adiabatic [555]. Adjacent [733, 589]. Admissibility [207]. Admittance [322]. Affine [557]. Albertoni [165]. Algebra [714]. Algebraic [311, 45, 614, 396, 169, 517]. Algebras [491, 516, 732, 511, 228, 607, 395]. Alloys [680]. Almost [87, 220]. Amplitude [583, 499, 634, 142, 143, 253, 479, 159, 160, 18, 480]. Amplitudes [476, 537, 662, 46, 526, 32, 429, 402, 425, 251, 691, 305, 500, 621]. Analog [665]. Analogs [175]. Analysis [711, 129, 365, 469, 560, 691, 462, 576, 684, 737]. Analytic [293, 213, 77, 88, 132, 288, 32, 199, 35, 49, 390, 201, 475, 586]. Analytical [79]. Analyticity [198, 152, 18]. Angle [182, 488]. Angular [186, 262, 679, 499, 281, 433, 689, 536, 548, 687, 734, 735, 275, 490, 282, 224, 26, 131]. Angular-Momentum [433, 689]. Anisotropic [560, 708]. Annihilation [406]. Antenna [239, 122, 171, 323, 553]. Antennas [322]. Anticommuting [249]. Antiequivalence [373]. Antiferromagnetism [641]. Antiorthogonal [65]. Antisymmetric [234, 655]. Antiunitary [42, 43]. Applicability [522]. Application [25, 662, 632, 294, 112, 722, 116, 206, 680, 691, 576, 684, 737, 41, 585, 553, 700]. Applications [110, 727, 129, 7, 98, 560, 461]. Approach [71, 486, 368, 512, 614, 701, 527, 70, 587, 695, 208, 348, 474, 136, 137, 240, 443, 36, 67]. Approximants [510]. Approximate [361, 487, 317, 531, 158, 56]. Approximation [539, 297, 581, 694, 588, 121]. Approximations [38, 107, 108]. Arbitrary [21, 63, 636, 182, 729, 542, 543, 181]. Arising [459, 617, 432]. Arrays [185]. Artificial [528]. Aspect [169]. Aspects [308, 23, 470]. Assemblies [8]. Associated [491, 683, 321, 445]. Asymptotes [497]. Asymptotic [60, 66, 599, 3, 404, 534, 268, 351, 354, 233, 409, 692, 500, 651, 285, 460, 133]. Asymptotically [277]. Atom [146, 161]. Atomic [306]. Atoms [185]. Attraction [417]. Averaging [309]. Avoiding [440, 644]. Axial [668]. Axially [626, 442]. Axiom [244]. Axiomatic [301, 379, 565]. Axioms [154, 448].

Baker [489, 263]. Bardeen [233]. Bargmann [637]. Bases [461]. Basic [579, 416]. Basis [262, 548, 687, 734, 735]. BCS [640]. be [596]. Behavior [499, 537, 123, 402, 219, 351, 428, 265, 277, 386, 498, 692, 563, 500]. Belonging [51]. Bergman [626]. Bessel [577, 651]. Bethe [536, 473, 474, 327, 621, 56]. Between [607, 544, 92, 57, 557, 454, 705, 246, 542, 43, 83]. beyond [400]. Bilinears [714]. Binary [370]. Bloch [267, 693, 317, 318]. Bocchieri [165]. Bodies [684, 525, 83]. Body [382, 297, 458, 231, 232, 584, 423, 285, 258, 306, 503]. Boltzmann [482, 140, 312, 216, 313, 38, 702, 13]. Bond [146, 643]. Born [16]. Bose [628, 491, 405, 382, 10, 11, 530, 605, 585, 73]. Boson [634, 66, 556, 280].

Bosons [567, 57, 623, 390, 407]. Bound [285]. Boundary [583, 123, 501, 75]. Boundary-Layer [123]. Bounded [636, 670]. Boundedness [681]. Bounds [205, 463, 730, 635, 636, 619, 90]. Box [422]. Branch [366]. Brownian [91, 493, 310, 639, 34, 105, 109]. Brownian-Motion [310].

Calculated [10]. Calculation [213, 52, 149, 488, 148, 125]. Calculations [531]. Calculus [551]. Canonical [572, 55, 405, 47, 401, 726, 471, 727, 622, 640, 204, 502, 591, 67]. Capacity [89]. Capture [627]. Carathéodory [25]. Carlo [62]. Case [640, 682, 473, 447, 403, 107, 108]. Cases [687, 735]. Casimir [341]. Cauchy [222, 320, 442]. Causal [202]. Causality [178, 568]. Caused [2]. Center [162, 417, 544]. Central [712]. Certain [459, 328, 402, 595, 651, 395]. Chains [194]. Changes [2]. Channel [451, 637, 226, 691, 80]. Character [151, 19]. Characteristic [28, 115, 279]. Characterization [45, 603]. Characters [658, 700]. Charge [111, 494, 705, 435, 633]. Charged [365, 426]. Chemical [459]. Circular [323]. Class [60, 402, 598, 541, 604, 617, 506, 256]. Classes [715]. Classical [30, 700, 716, 164, 549, 329, 388, 507, 309, 454, 535, 165, 14, 343, 423, 247, 701]. Classification [726, 98, 82]. Clebsch [738, 673]. Cluster [128, 127, 191, 444, 645, 503]. Coaxial [322]. Coefficient [288]. Coefficients [110, 635, 215, 216, 314, 513, 240, 443, 409, 693, 633]. Coherence [252]. Collisionless [13]. Collisions [72, 145, 297]. Combinatorial [486, 522, 23, 470]. Comments [661]. Commutation [55, 92, 405, 401]. Commutative [156]. Commutativity [475]. Commutator [27, 50]. Commutators [248]. Compact [732, 338]. Comparison [730, 494, 146, 410]. Complete [563]. Completion [132]. Complex [679, 713, 476, 451, 330, 188, 189, 190, 411, 281, 429, 193, 65, 412, 275, 490, 18, 364, 553, 295]. Complex-Order [553]. Component [20, 462]. Components [306]. Composite [458, 575, 574]. Compound [100]. Compressibility [455]. Computation [62, 478]. Computations [450]. Computer [483, 410]. Concept [675]. Concepts [416]. Concerning [564, 28, 698, 331]. Condensate [347]. Condition [596, 327, 133]. Conditions [572, 583, 501, 415, 207, 138, 24, 212, 620]. Conducting [697]. Conduction [15, 167, 130]. Conductivity [678, 215, 574]. Conductor [217]. Conductors [512]. Configuration [678]. Configurations [238, 181, 364]. Conformal [592]. Congruences [711]. Conjecture [260, 23, 470, 292]. Conjugation [725]. Connection [92, 557, 607, 675]. Conservation [557, 593, 715]. Consistency [47]. Consistent [308]. Constant [60]. Constraints [307]. Construction [377, 732, 492, 141, 180, 707, 736, 545]. Containing [278, 651]. Continuant [293]. Continuation [88, 475]. Continued [148]. Continuous [453, 454, 535, 617, 291]. Continuous-Representation [453, 454, 535, 617]. Continuum [210]. contra [507]. Contraction [64, 150]. Convergence [16, 599, 203, 556, 613, 485]. Convergent [595, 221]. Convex [261]. Cooper [233]. Coordinate [44, 506, 138]. Coordinates [416]. Core [469, 307].

Cores [567, 554]. Corrections [288, 682]. Correlated [306]. Correlation [690, 624, 520, 449, 22, 633]. Correlations [587, 358, 630]. Coulomb [599, 264, 162, 704, 581, 659, 210, 139, 391, 389, 570, 115, 561, 369, 706, 78]. Counting [68, 243]. Coupled [526, 501, 393, 410, 8, 265, 664, 667]. Covariance [299, 420]. Covariant [471, 271]. Creation [494]. Criteria [394]. Criterion [303, 151]. Critical [128, 624, 519, 643, 120]. Cross [449, 182, 331]. Cross-Correlation [449]. Crossed [691]. Crossing [691]. Crystal [195, 193, 263]. Crystals [103, 693, 2, 104]. Cube [376]. Cubic [34, 91, 105, 192]. Current [217]. Curvature [60]. Curves [134, 497]. Cusps [134, 430]. Cutoff [203]. Cuts [366]. Cycles [8]. Cyclic [213, 629, 52]. Cyclotron [218]. Cylinder [678, 575, 426]. Cylinders [182, 538].

Damped [529]. Decays [359]. Decomposition [29]. Deductions [331]. Deep [610]. Defining [168]. Definition [248, 100]. Deformations [427]. Deformed [15]. Degree [104]. Dense [345]. Density [539, 716, 39, 155, 272, 513, 521, 432, 372]. Dependence [104]. Dependent [482, 357, 294, 112, 272]. Derivation [385, 17, 39, 687, 735, 533, 207, 289, 590]. Derived [237]. Describing [405]. Description [59, 493, 9]. Determinant [462]. Determinantal [514, 657, 558]. Determinants [234, 168]. Determination [86]. Development [340, 582, 503]. Diagonalization [450]. Diagonalization-Some [450]. Diagram [13]. Diagrammatic [518]. Diagrams [563]. Diatomic [72]. Dielectric [423]. Different [92]. Differential [697, 3, 582, 416, 693, 562, 551, 445, 38, 396, 169]. Differential-Operator [38]. Differentiation [418]. Diffraction [611, 538, 333, 740]. Diffusion [588]. Dilute [680, 642, 266]. Dimension [57]. Dimensional [666, 205, 76, 249, 298, 264, 433, 689, 527, 349, 665, 124, 139, 391, 623, 373, 584, 606, 358, 290, 719, 720, 261, 618, 543, 82, 407, 724, 2, 107, 108]. Dimensions [377, 481, 339, 304, 643, 181]. Dimer [356]. Dimers [316]. Dipole [469, 239, 122, 171, 322]. Dirac [712, 179, 502, 447, 714, 597, 561, 276]. Dirac-Type [179]. Direct [29]. Disconnected [632]. Discontinuities [46]. Discrete [687, 735]. Discussion [519, 349, 350]. Dislocations [2]. Disordered [124, 103, 194, 104]. Dispersion [474, 677]. Displacement [183]. Distance [174, 542, 544]. Distinct [467]. Distinction [43]. Distortion [417]. Distribution [572, 164, 10, 11, 351, 378, 623, 350]. Distributions [21, 63, 321, 222, 320, 88, 111, 479, 603, 719, 720, 255, 257, 480]. Divergence [415]. Domains [274, 273]. Double [238, 474, 27, 50]. Double-Tensor [238]. Down [112]. Dressed [359]. Dressed-Particle [359]. Driven [322]. Dual [705]. Duality [606, 686]. Due [422, 417]. Dyadic [711, 670]. Dynamic [265]. Dynamical [155, 272, 605]. Dynamics [549, 329, 388, 566, 454, 74, 196, 591, 34, 91, 105]. Dyson [132, 260, 292].

Eckart [413]. Edge [58]. Edge-of-the-Wedge [58]. Effect [342, 469, 729]. Effective [699, 117, 5]. Effective-Range [117]. Effects [294, 61]. Eigenfunction [389, 115]. Eigenfunctions [282, 363, 600]. Eigenvalue [716, 147, 286, 441]. Eigenvalues [229, 730, 539, 310, 90, 149, 184, 148, 214]. Eigenvectors [184, 148]. Einstein [60, 98, 99, 438, 136, 137, 209, 114]. Elastic [205, 537, 374, 573, 427, 100]. Electric [391, 705]. Electrical [215, 341, 15, 167, 130]. Electrodynamics [550, 299, 174, 221, 668, 1]. Electrolytic [342]. Electromagnetic [269, 176, 183, 534, 575, 663, 573, 718, 593, 710, 729, 715, 94, 669]. Electromagnetism [271]. Electron [21, 63, 40, 397, 215, 290, 600, 364]. Electrons [704, 238, 466, 602]. Electrostatic [512, 364]. Elementary [661, 632, 48]. Elementary-Particle [632]. Elements [205, 636, 578, 478, 707, 736, 421]. Ellipsoidal [111]. Emission [550]. Empty [136, 97, 277, 278, 436]. Empty-Space [436]. Energy [463, 482, 374, 40, 397, 368, 188, 189, 190, 411, 106, 483, 404, 402, 656, 260, 112, 422, 604, 603, 378, 332, 259, 227, 254, 412, 261, 180, 559, 390, 386, 498, 563, 427, 477, 158, 331, 214]. Energy-Dependent [482, 112]. Energy-Level [604, 603]. Ensemble [30, 309, 480]. Ensembles [311, 604]. Entropy [517, 518]. Equation [493, 683, 666, 446, 646, 685, 482, 638, 361, 514, 375, 140, 495, 54, 536, 442, 520, 90, 672, 665, 313, 657, 731, 473, 474, 552, 648, 159, 160, 597, 276, 327, 670, 237, 38, 702, 652, 170, 326, 392, 37, 353, 12, 13, 56, 590, 586, 26, 131, 121]. Equations [380, 471, 494, 697, 510, 661, 501, 3, 592, 179, 135, 449, 541, 344, 582, 509, 709, 85, 22, 136, 137, 339, 650, 693, 562, 415, 521, 466, 602, 445, 209, 317, 289, 667, 396, 169, 567]. Equilibrium [71, 646, 287, 519, 520, 70, 587, 349, 231, 559, 36, 350]. Equipartition [106]. Equivalence [373, 645]. Equivalent [238]. Ergodic [393]. Ergodicity [483, 8, 24, 212]. Errata [734, 735, 443, 490, 91, 444, 737, 628, 131]. Erratum [269, 397, 738, 242, 296, 739, 196, 736, 150, 243, 740]. Estimating [229]. Estimation [315]. Eulerian [208]. Evaluation [459, 312, 647, 739, 95, 96, 487]. Evolution [195]. Exact [572, 666, 264, 355, 721, 139, 391, 308, 126, 473, 282, 561, 652, 643, 125, 291, 120, 22]. Exactly [464, 584]. Examples [204]. Excitation [628, 530, 235]. Excluded [69, 242, 68, 243]. Existence [638, 495, 593, 339]. Expansion [16, 546, 162, 173, 721, 95, 96, 332, 115, 731, 233, 221, 542, 544, 651, 41, 392]. Expansions [577, 69, 242, 203, 613, 22, 485, 595, 445, 543]. Expectation [324, 325, 400]. Explicit [725, 377, 545, 689]. Expression [5]. Extended [9]. Extension [569]. Extensive [656]. External [515]. External-Mass [515]. Extremely [73].

far [646]. Faxén [590]. Fermi [739, 647, 416, 266]. Fermion [464, 235]. Fermions [57]. Fermis [717]. Ferrite [663]. Ferro [641]. Ferro- [641]. Ferromagnetics [642]. Ferromagnetism [370, 23, 470]. Feynman [470, 328, 662, 46, 298, 251, 552, 692, 487, 270, 23]. Field [567, 380, 55, 77, 491, 516, 324, 325, 361, 195, 269, 176, 66, 330, 153, 596, 534, 294, 173, 400, 200, 614, 469, 157, 532, 244, 511, 449, 84, 113, 391, 709, 308, 199, 492, 715, 719, 720, 653, 136, 137, 415, 360, 597, 565, 675, 114, 27, 50, 201, 175, 154, 448, 31, 67, 247, 133, 585, 669]. Field-Theory [400, 175]. Fields [92, 76, 249, 298, 274, 173, 135, 502, 655, 151, 715, 246, 94, 317, 675, 93, 545, 280, 724, 158, 247]. Final [359]. Final-State [359]. Finite [382, 632, 556, 28, 710, 232, 373]. First [683, 221, 445]. First-Order [445]. First-Quantized [221]. Five [172]. Five-Point [172]. Fixed [180]. Flat [277, 506]. Flat-Space [506]. Flatter [437]. Flow [625, 217]. Fluctuation [125]. Fluctuations [627, 346]. Fluid [610, 625]. Fluids [164, 624, 102, 509, 336, 485]. Folk [173]. Forced [170]. Forces [264, 520, 456, 139, 391, 246]. Form [471, 404, 389, 354, 409, 42]. Formal [327, 545, 81, 12]. Formalism [567, 55, 324, 325, 379, 674, 101, 286, 37, 255, 257]. Formalisms [620]. Forms [295]. Formula [132, 385, 17, 606, 263]. Formulas [662, 424]. Formulation [140, 329, 388, 517, 518, 458, 370, 174, 709, 591]. Formulations [287, 271]. Forward [537]. Foundations [197, 161]. Four [431, 99, 82]. Four-Dimensional [82]. Four-Parameter [99]. Four-Point [431]. Fourier [88, 635]. Fourth [652, 18]. Fourth-Power [652]. Fraction [148]. Fredholm [514, 657]. Free [380, 491, 405, 516, 471, 472, 596, 656, 95, 96, 151, 261, 93, 255]. Free-Field [380]. Free-Particle [471]. Free-Space [255]. Frequency [124, 166]. Front [59]. Fugacity [485, 508]. Function [77, 683, 198, 514, 172, 315, 404, 10, 11, 431, 520, 659, 210, 351, 389, 570, 116, 115, 206, 354, 617, 233, 252, 423, 192, 651, 706, 285, 398, 395, 475, 78]. Functional [264, 309, 7, 535, 695, 233, 551, 487]. Functionals [378]. Functions [293, 213, 77, 86, 374, 577, 406, 624, 234, 301, 152, 523, 596, 581, 721, 425, 654, 273, 126, 51, 22, 612, 141, 35, 49, 562, 207, 390, 319, 561, 542, 543, 31, 306, 376, 350, 266, 318]. Fundamental [496, 688]. Furry [581].

Galilei [419]. Galvanomagnetic [294, 61]. Gas [405, 355, 39, 465, 648, 407, 628, 530, 605, 664, 585]. Gases [345, 216, 408, 513, 485, 343, 424]. Gauge [299, 668, 1]. Gell [385]. Gell-Mann [385]. General [567, 71, 140, 330, 420, 70, 473, 207, 36, 170, 525, 256, 658, 47, 215]. General-Mass [473]. Generalization [58, 671, 722, 436, 383, 542, 41]. Generalizations [494, 612]. Generalized [77, 635, 596, 555, 626, 457, 449, 579, 437, 731, 521, 362, 363, 423, 289, 392, 298, 454]. Generating [721]. Generation [708]. Generators [223]. Geodesic [278]. Geometry [482, 696, 560, 416, 686, 83]. Ghosts [414]. Given [66, 707, 736]. Global [25, 399, 489]. Goldberg [383]. Gordan [738, 673]. Gordon [222, 320, 731, 326]. Graphs [662, 129, 23, 470]. Gravitational [135, 84, 502, 417, 240, 443, 415, 597, 279, 158, 247]. Graviton [676]. Gravity [625]. Green [86, 321, 222, 320, 514, 581, 659, 283, 389, 570, 115, 390, 706, 670, 31, 78]. Ground [623]. Group [380, 679, 713, 110, 377, 434, 569, 98, 592, 112, 532, 481, 616, 65, 419, 29, 223, 548, 687, 734, 735, 302, 568]. Group-Theoretical [112, 532]. Groups [658, 700, 505, 725, 726, 377, 76, 249, 298, 484, 311, 338, 413, 99, 302, 51, 461, 607, 64, 19, 150].

Half [733, 589]. Half-Spaces [733, 589]. Hamiltonian [55, 549, 507, 690, 174, 268, 620, 306]. Hamiltonians [212]. Hankel [425]. Hard [10, 11, 628, 530, 605, 307, 585, 73, 266]. Hard-Sphere [628, 530, 605, 585]. Harmonic [546, 335, 8]. Harmonic-Type [546]. Harmonically [170]. Harmonics [140, 543]. Hausdorff [489, 263]. Heavy [300]. Heisenberg [690, 274, 370, 641, 642]. Helicity [691]. Hermitian [149]. High [402, 227, 719, 720, 386, 498, 563, 192, 477, 331]. High-Energy [402, 227, 386, 498, 563, 477, 331]. High-Field [719, 720]. High-Temperature [192]. Higher [107, 108]. Higher-Order [107, 108]. Hilbert [17, 631]. Hole [682]. Holomorphy [273]. Holonomy [98, 99]. Homomorphically [302]. Hose [677]. Hove [212]. Hulthén [4]. Hulthén-Type [4]. Hydrodynamical [161]. Hydrodynamics [347, 528]. Hydrogenlike [319]. Hyperbolic [582]. Hypergeometric [562]. Hypersurface [278].

Icosahedron [89]. Identical [211]. If [717]. II [567, 262, 700, 505, 325, 388, 518, 249, 189, 264, 682, 379, 216, 72, 689, 143, 70, 410, 272, 226, 644, 454, 387, 232, 96, 206, 571, 719, 496, 160, 609, 498, 105, 448, 306, 257, 350, 108]. III [71, 725, 298, 190, 314, 145, 253, 519, 535, 391, 22, 720, 688]. Imaginary [28]. Imbedded [697]. Imbedding [33]. Impedance [649, 202]. Impenetrable [567, 57, 623, 407]. Imperfect [39, 408]. Implies [568]. Impurities [540, 571]. Impurity [628, 530, 605]. Incomplete [432]. Inequalities [163, 690, 90, 509]. Infinite [405, 678, 76, 249, 298, 640, 168, 231]. Infinite-Dimensional [76, 249, 298]. Infinitely [610, 322]. Infinitesimal [98, 223]. Infinitesimal-Holonomy [98]. Infinity [60]. Influence [587]. Inhomogeneity [719, 720]. Inhomogeneous [646, 616, 29, 223, 548, 687, 734, 735, 719, 720, 160, 252, 71]. Initial [696, 587, 438, 279]. Initial-Value [696, 438]. Input [322]. Instability [677]. Instance [606]. Insulators [667]. Integrable [478]. Integral [683, 510, 429, 523, 626, 672, 509, 22, 254, 305, 496, 688, 78]. Integrals [328, 577, 647, 739, 552, 692, 561, 487, 651, 318]. Integration [264, 284, 7, 562, 169, 26, 131]. Integrodifferential [650]. Intensive [580]. Interacting [556, 296, 187, 502, 259, 390, 247, 660, 70, 36]. Interaction [549, 507, 599, 653, 478]. Interactions [634, 193, 598, 359, 246, 364]. Interatomic [193]. Interest [698]. Interface [729]. Intermediate [634]. Intermediate-Vector-Boson [634]. Interpretation [114]. Introduction [284]. Invariance [76, 249, 298, 555, 592, 84, 686, 326]. Invariant [33, 674, 481, 473]. Invariants [377, 732, 427, 175, 18]. Inverse [334, 185, 52, 698, 652]. Inverse-Scattering [698]. Inversion [629]. Involving [494, 543]. Irreducibility [244]. Irreducible [658, 700, 546, 457, 373, 461, 224, 382]. Irregular [424]. Irregularities [729]. Irregularly [512]. Irreversibility [8]. Irreversible [371]. Ising [242, 444, 470, 243, 486, 69, 62, 337, 357, 606, 265, 358, 409, 191, 192, 642, 23, 630, 68, 243]. Ising-Model [409, 630]. Isomorphic [129]. Isoperimetric [163, 90]. Isotherm [355]. Isotopically [103, 104]. Isotropic [374, 15, 61]. Iterated [222]. Iterative [685]. IV [726, 411, 297, 520, 286, 617].

J [740, 184]. Jost [132, 612].

Kahan [722]. Kemmer [54, 607]. Kernel [514, 370, 560]. Kind [683]. Kinematics [622, 384]. Kinetic [446, 646, 345, 701, 465, 466, 602, 664, 723]. Klein [222, 320, 731, 326]. Kramers [606].

Ladder [563]. Lagrange [41, 175]. Lagrangian [709, 208, 620, 133]. Lamb [95, 96]. Lamm [590]. Landau [134, 497]. Lanford [283]. Langevin [493]. Laplace [442, 542]. Large [499, 577, 422, 417, 468]. Lattice [491, 40, 397, 62, 335, 355, 219, 522, 456, 124, 290, 485, 166, 192, 630, 467]. Lattices [193, 34, 91, 105, 316]. Law [593]. Laws [557, 715]. Layer [123]. Legendre [683]. Lehmann [132, 619]. Lemma [569]. Lemmas [173]. Level [604, 603]. Levels [463, 188, 189, 190, 411, 260, 259, 412, 290]. Lie [150, 505, 725, 726, 377, 413, 732, 228, 64, 395, 396]. Light [177, 161, 83]. Limit [640]. Limitation [455]. Limiting [159, 160]. Line [335, 122]. Linear [418, 697, 510, 641, 655, 428, 194, 38, 396, 169]. Linearized [21, 63]. Lines [322]. Liouville [344, 37, 12]. Liquid [519, 520, 349, 350]. Liquids [408]. Local [679, 516, 25, 532, 511, 475]. Local-Field [532]. Localization [113, 504]. Locally [236]. Loinger [165]. London [339]. Long [520, 219, 456, 332, 648, 117, 322]. Long-Range [520, 456, 332, 648, 117]. Long-Time [219]. Loop [323]. Lorentz [734, 713, 727, 674, 84, 616, 65, 29, 223, 548, 687, 735, 302, 271, 424, 568]. Lorentzian [93]. Lorenz [424]. Low [368, 332, 513, 317, 73]. Low-Density [513]. Low-Energy [368, 332]. Lower [463, 730, 635, 636, 619, 90]. Lyapunov [394].

Magnetic [678, 469, 317]. Magnetically [680, 667]. Magnetization [690, 358]. Magnetomechanical [667]. Magnetoplasmas [670]. Magnetoresistance [719, 720]. Magnus [489]. Mandelstam [300, 430, 681, 6]. Manifolds [438]. Mann [385]. Many [679, 451, 637, 404, 458, 231, 232, 464, 141, 80, 423, 31, 306]. Many-Body [458, 231, 232, 423, 306]. Many-Channel [451, 637, 80]. Many-Fermion [464]. Many-Particle [679, 404, 141, 31]. Mappings [155, 272]. Markovian [572]. Mass [320, 111, 385, 592, 251, 616, 223, 687, 735, 417, 515, 473]. Mass-Zero [592]. Masses [431, 468]. Massless [715]. Master [521, 237, 289, 392, 353]. Material [83]. Materials [374, 15, 61]. Math [740]. Mathematical [33, 469, 331]. Matrices [293, 213, 186, 262, 539, 716, 629, 691, 156, 432, 295]. Matrix [205, 79, 451, 147, 636, 310, 178, 566, 185, 173, 118, 387, 479, 28, 52, 578, 478, 80, 275, 490, 447, 521, 149, 450, 698, 707, 736, 184, 392, 421, 306, 372]. Matter [247, 660]. Maxwell [494, 137, 339, 209, 271]. Maxwell-Einstein [137]. Maxwellian [236]. Mayer [39, 508]. Means [233]. Measurements [432]. Mechanical [418, 312, 225, 226, 207, 360]. Mechanics [59, 524, 564, 147, 546, 507, 703, 287, 352, 311, 264, 197, 631, 204, 155, 272, 139, 391, 419, 85, 101, 165, 617, 680, 24, 212, 362, 363]. Media [697, 718]. Medium [649, 573, 574, 160, 252]. Meson [203, 619]. Method [264, 62, 371, 309, 10, 626, 296, 187, 522, 721, 722, 408, 14, 579, 168, 492, 233, 240, 443, 441, 558, 121, 19]. Methods [730, 569, 4, 576, 684, 737, 702]. Metric [713, 417, 436]. Metrics [278]. Microns [717]. Microscopic [348]. Microwave [649]. Minimum [261, 4]. Model [69, 242, 310, 640, 337, 357, 370, 400, 359, 349, 641, 464, 606, 358, 409, 360, 34, 91, 105, 191, 444, 23, 470, 630, 68, 243, 724, 125, 161, 660, 291]. Models [661, 532, 231, 232]. Moderation [695]. Modes [556, 193, 168, 36, 194]. Modification [117]. Modified [493, 300, 694, 317]. Moduli [707, 736]. Molecular [463, 72, 145, 297]. Molecules [72, 145]. Momenta [679, 281, 275, 490]. Moments [378, 237]. Momentum [186, 262, 77, 499, 198, 284, 433, 689, 274, 659, 623, 548, 687, 734, 735, 282, 362, 363, 224]. Momentum-Space [198]. Monoenergetic [696]. Monte [62]. Motion [567, 493, 310, 557, 639, 14, 34, 91, 105, 109, 684, 529, 525, 258]. Moving [75, 718]. MR [740]. Multi [320, 226]. Multi-Channel [226]. Multi-Mass [320]. Multichannel [674]. Multiple [682, 112, 654, 710, 627, 588, 181, 119, 107, 108]. Multiple-Valued [654]. Multiplets [571]. Multiplication [88]. Multiplicative [565]. Multipole [144]. Mutual [252].

n [481, 304]. Nature [154, 448]. Navier [665]. Near [266]. Nearest [690]. Nearest-Neighbor [690]. Nearly [268, 230, 297]. Necessary [596]. Neighbor [690]. Nested [631]. Network [129]. Networks [341, 699]. Neumann [491, 516, 511]. Neutron [495, 163, 722, 378, 560, 733, 589, 121]. Neutron-Transport [121]. Neutrons [696, 112]. Newman [437]. No [549, 26, 131]. No-Interaction [549]. Nodal [22]. Non [217]. Non-Rectilinear [217]. Noncircular [167]. Noncrossing [192]. Nonequilibrium [340, 232, 103, 237, 73, 104]. Nonexistence [60]. Nonlinear [361, 661, 106, 483, 393, 410, 74, 196, 541, 582, 168, 559, 53, 545, 125, 346, 291]. Nonlinearity [731]. Nonlocal [598, 650, 67, 585]. Nonlocalizable [675]. Nonorthogonality [363]. Nonrelativistic [405, 524, 659, 419, 389, 570, 653, 275, 490, 285, 460, 78]. Nonrenormalizable [675]. Nonsymmorphic [19]. Nonuniform [336].

Nonunitary [484]. Nonzero [687, 735]. Norm [592]. Normal [352, 517, 518, 173, 416, 36, 42, 295]. Normalization [327]. Note [401, 615, 375, 647, 739, 165, 230, 156, 407, 280, 489, 169]. Nuclear [478]. Nucleon [118, 387]. Nucleon-Nucleon [118, 387]. Null [506, 94]. Number [382, 556, 440, 644, 290, 467]. Numerical [315, 721, 450, 528, 26, 131].

Obeying [216]. Object [611]. Objects [697]. Observable [468]. Observables [516, 703, 352, 511, 580]. Obtain [19]. Octet [578]. Okubo [385]. One [666, 205, 264, 162, 57, 527, 349, 665, 124, 139, 391, 623, 584, 733, 589, 290, 407]. One- [162]. One-Dimensional [666, 205, 264, 527, 349, 665, 124, 139, 391, 623, 584, 290, 407]. One-Speed [733, 589]. Onsager [341]. Operation [725, 45]. Operational [576, 684, 737]. Operations [615]. Operator [567, 321, 222, 320, 472, 636, 626, 433, 689, 17, 244, 101, 286, 578, 116, 206, 254, 343, 38, 202, 723, 214]. Operators [418, 567, 524, 726, 546, 406, 155, 272, 457, 238, 481, 373, 492, 478, 595, 362, 551, 600, 201, 42, 43, 585]. Optical [540, 571]. Orbital [433, 689]. Order [577, 62, 400, 603, 445, 18, 553, 107, 108, 104]. Ordered [173]. Ordering [259]. Ordinary [169]. Oriented [182]. Orthochronous [223]. Orthogonal [65, 278]. Orthogonality [319, 307]. Oscillating [195]. Oscillations [21, 63]. Oscillator [483, 457, 109, 529]. Oscillators [393, 410, 8]. Other [163, 573, 715, 268]. Overlap [185, 52].

Padé [510]. Pair [10, 11, 520, 445]. Pais [633]. Paper [283]. Parameter [418, 99]. Parameters [62, 428]. Partial [583, 697, 32, 365, 582, 691, 693, 562, 379]. Partial-Wave [583, 691]. Particle [679, 676, 549, 471, 622, 507, 632, 566, 712, 404, 728, 17, 359, 695, 384, 22, 141, 478, 435, 343, 407, 31, 100, 403, 33]. Particles [472, 661, 48, 211, 296, 187, 458, 536, 70, 365, 135, 300, 653]. Partition [315, 233, 192, 395]. Parts [28, 390]. Passive [699]. Pathological [698]. Paths [23, 470, 83]. Penetrable [333, 740]. Peratization [609]. Percolation [128, 127, 146, 643]. Percus [586]. Periodic [406, 649, 185, 268, 693, 230, 602]. Periodicity [87, 220]. Perpendicular [337]. Perturbation [172, 512, 284, 203, 301, 379, 402, 303, 431, 541, 286, 308, 731, 228, 305, 496, 688, 386, 498, 563, 595, 230, 221, 558, 702, 460, 601, 393]. Perturbation-Theoretical [496, 688]. Perturbational [170]. Petrov [98]. Pfaffian [522]. Phase [17, 527, 356, 332, 591, 180, 694, 531, 100]. Phase-Space [591]. Phenomena [701]. Phenomenological [533, 43]. Phonon [215, 709]. Phonons [466]. Phys [740]. Physical [179, 254, 41]. Physics [459, 33, 7, 263]. Picture [359]. Planar [403]. Plane [482, 355, 258]. Plasma [21, 63, 446, 646, 140, 236, 723]. Plasmas [394, 719, 720, 576, 737]. Poincaré [8, 175]. Point [40, 397, 172, 431, 273, 407, 398]. Point-Particle [407]. Points [666, 3, 542, 544]. Poisson [606]. Pol [170]. Polarization [524]. Poles [366, 526, 554, 414, 425, 275, 490, 477]. Polygonal [538].

Polygons [192]. Polynomial [681, 121]. Polynomials [716]. Position [524, 472]. Positive [648]. Positronium [236]. Possibilities [157]. Possible [705, 399]. Postulates [200, 453]. Potential [79, 367, 366, 368, 32, 712, 281, 10, 626, 422, 6, 227, 648, 117, 447, 452, 652, 477]. Potentials [637, 554, 162, 332, 180, 609, 601]. Power [652]. Powers [542, 544]. Presence [718, 117]. Preservation [572]. Pressure [39]. Primitive [273]. Principal [168, 81]. Principle [517, 518, 248, 420, 211]. Principles [533, 159, 160, 207]. Probabilities [128, 643, 353]. Probability [164]. Problem [20, 696, 222, 320, 132, 486, 123, 69, 242, 512, 334, 728, 458, 442, 393, 225, 226, 75, 438, 241, 286, 208, 141, 698, 279, 68, 243, 306, 263, 503, 553, 120]. Problems [147, 128, 127, 522, 584, 441, 258, 41, 643, 256]. Procedure [328, 229, 541, 594]. Processes [572, 33, 371, 146, 103, 627, 237, 104]. Produced [307]. Product [645]. Production [476, 305]. Products [186, 262, 29, 201, 81]. Projection [286, 282]. Prolate [469]. Proof [25, 656, 260, 579, 261, 292]. Propagator [414, 95, 96]. Propagators [1]. Proper [404, 616, 223]. Properties [380, 79, 198, 727, 662, 66, 172, 288, 32, 129, 598, 534, 468, 496, 688, 97, 35, 49, 390, 224, 547, 285, 605, 201, 73, 202, 372, 1, 679]. Property [4, 319, 645]. Proximity [417]. Pseudo [10]. Pseudo-Potential [10]. Pure [211]. Purely [533].

Q [420]. Quadratic [427]. Quantal [87, 220]. Quantities [449]. Quantization [666, 76, 249, 298, 535, 594, 270, 53, 668]. Quantized [653, 221, 93, 247]. Quantum [59, 418, 55, 77, 491, 324, 325, 550, 299, 564, 546, 329, 388, 703, 288, 312, 311, 381, 334, 197, 315, 216, 7, 296, 187, 400, 631, 200, 614, 204, 155, 272, 225, 226, 244, 511, 454, 113, 419, 313, 199, 617, 492, 207, 360, 24, 212, 565, 362, 363, 602, 109, 545, 702, 27, 50, 580, 154, 448, 31, 1, 314]. Quantum-Mechanical [418, 312, 225, 226, 207]. Quaternion [197]. Quaternionic [338]. Quenched [680].

R [178]. Radial [351, 35, 49, 319, 561, 276]. Radiation [550, 718, 240, 443, 218, 144, 660]. Rainich [20]. Random [539, 716, 697, 310, 219, 573, 479, 378, 695, 654, 710, 290, 450, 255, 257, 467]. Randomly [428, 182, 642]. Range [520, 456, 332, 648, 117, 5]. Rate [237]. Ray [182, 488]. Rays [278]. Real [713, 451, 44, 28, 354]. Realizations [223]. Reciprocal [341]. Reciprocity [564]. Rectangular [316]. Rectilinear [217]. Reduced [479, 480, 372]. Reduction [47, 569, 728, 244, 421]. Regge [537, 366, 554, 414, 497, 477]. Reggeization [691]. Region [624, 519]. Regions [437]. Regularity [447]. Regularization [199]. Related [132, 662, 302, 192]. Relating [716, 690]. Relation [454, 677]. Relations [55, 92, 405, 401, 341]. Relationship [57, 705]. Relativistic [59, 524, 549, 472, 622, 507, 703, 566, 704, 465, 384, 199, 468, 452, 276, 369, 529, 326, 154, 448]. Relativity [330, 525, 47]. Relaxation [572, 237]. Remainder [508]. Remarkable [434, 607]. Remarks [564, 486, 681, 511, 28, 474, 698, 148]. Renormalization [634, 517, 518, 247]. Representation [451, 406, 484, 434, 640, 300, 430, 681, 453, 454, 535, 655, 6, 65, 617, 258, 724, 78]. Representations [679, 658, 700, 405, 382, 505, 725, 726, 377, 44, 338, 429, 523, 569, 689, 457, 616, 29, 548, 687, 734, 735, 570, 461, 305, 496, 688, 547]. Repulsive [601]. Resistance [699]. Resolvent [37]. Respect [306]. Response [239, 171]. Restricted [302]. Restrictions [294]. Results [579]. Retarded [77, 301]. Review [355]. Ribbon [333, 740]. Riccati [375]. Rideau [722]. Riemann [54]. Riemannian [439, 686, 82]. Rigid [297, 684]. Rigorous [17]. Ring [312]. Rings [244]. Rod [167]. Role [133]. Root [326]. Roots [28]. Rosen [438]. Rotation [679]. Rotational [282]. Rotations [44]. Roussopoulos [722]. Rules [666, 479, 331, 318].

Sachs [383]. Salpeter [536, 473, 474, 327, 621, 56]. Saturated [667]. Scalar [516, 330, 682, 135, 665, 653, 107, 108]. Scatterer [255, 257]. Scatterers [710]. Scattering [205, 79, 499, 367, 638, 537, 610, 608, 366, 697, 368, 526, 334, 682, 32, 712, 281, 429, 402, 704, 17, 142, 143, 253, 112, 359, 365, 575, 225, 226, 674, 663, 425, 672, 6, 118, 387, 332, 227, 116, 206, 710, 254, 182, 729, 159, 160, 447, 452, 698, 343, 369, 488, 588, 500, 280, 531, 621, 81, 477, 18, 100, 256, 181, 255, 257, 161, 119, 403, 107, 108]. Schemes [399]. Schrieffer [233]. Schrödinger [666, 638, 501, 142, 143, 253, 90, 552, 159, 500, 114, 652]. Schur [569]. Schwartz [88]. Schwarzschild [437, 436]. Schwinger [298]. Second [288, 245]. Section [182]. Sections [331]. Segal [594]. Selection [318]. Self [404, 440, 644, 308, 390]. Self-Avoiding [440, 644]. Self-Consistent [308]. Self-Energy [404, 390]. Sellmeier [424]. Semi [129]. Semicircle [539]. Semiconductors [719, 720]. Semisimple [505, 725, 726, 377, 732]. Separation [536]. Series [673, 738, 69, 242, 39, 409, 508]. Set [265]. Sets [373]. Several [456]. Shape [335, 729, 261]. Shaped [512]. Sharing [483, 559]. Sheet [254]. Shells [427]. Shift [17, 95, 96, 100]. Shifts [422, 332, 180, 694, 531]. Shmushkevich [579]. Sign [479]. Simple [658, 700, 624, 413, 465, 718, 223, 34, 91, 105, 192, 395, 590]. Single [17, 462]. Single-Determinant [462]. Singlet [634]. Singlets [540]. Singular [672, 116, 206, 609, 445, 621]. Singularities [476, 46, 3, 245, 429, 303, 431, 251, 582, 497, 81]. Singularity [515]. Site [643]. Sites [467]. Sitter [434]. Size [128, 127, 710]. Slab [696, 560, 120]. Slater [234]. Slowing [112]. Slowing-Down [112]. Small [417, 182, 729, 731, 488]. Small-Angle [488]. Smooth [611]. Solid [546, 680]. Solids [608, 540, 571]. Soluble [464, 584]. Solution [685, 361, 696, 514, 510, 442, 672, 85, 241, 657, 339, 327, 652, 170, 12, 13, 396, 590, 586, 291, 120, 286]. Solutions [638, 3, 236, 179, 342, 410, 449, 665, 268, 582, 680, 473, 317, 56]. Solvable [598]. Solving [441]. Some [110, 662, 172, 127, 647, 739, 98, 400, 179, 204, 28, 416, 461, 97, 409, 390, 450, 698, 561, 547, 605, 68, 243, 73, 202, 331, 372, 308]. Sommerfeld [553]. Soret [342]. Source [66]. Space [77, 198, 60, 54, 153, 406, 711, 640, 284, 17, 274, 659, 591, 437, 136, 97, 278, 506, 436, 255, 257, 19]. Space-Time [711, 97]. Spaces [60, 7, 98, 631, 557, 99, 439, 373, 617, 733, 589, 277, 618, 82]. Spacetime [48]. Spacing [603]. Special [353]. Spectra [604, 540, 571]. Spectral [560, 276]. Spectrum [632, 124, 254, 166, 628, 530, 125, 235]. Speed [733, 589]. Sphere [697, 10, 11, 628, 530, 605, 585]. Spheres [73, 266]. Spherical [140, 546, 297, 543, 262]. Spheroid [426]. Spheroidal [469]. Spin [92, 250, 269, 176, 676, 727, 690, 656, 536, 619, 259, 687, 735, 240, 443, 692, 462, 600, 630, 669]. Spin- [536]. Spin-2 [676]. Spin-Component [462]. Spin-Statistics [250]. Spin-Zero [619]. Spinless [365]. Spinor [655]. Spinors [20, 304]. Spins [265]. Spontaneous [358]. Square [326]. Square-Root [326]. Stability [394, 625]. State [359, 520, 623, 648]. States [679, 382, 211]. Static [60, 339]. Stationary [517, 518, 5]. Statistical [287, 352, 501, 188, 189, 190, 411, 264, 314, 260, 232, 336, 604, 428, 139, 391, 85, 101, 165, 533, 412, 680, 34, 91, 105, 161]. Statistics [250, 269, 176, 312, 355, 315, 216, 296, 187, 357, 356, 313, 669, 391]. Stimulated [550]. Stochastic [697, 74, 196, 231, 232, 344]. Stokes [665]. Straight [217]. Strain [374, 427]. Stratified [625]. Stress [158]. Stretched [130]. Strict [113, 504]. String [291]. Strip [241]. Stroboscopic [541]. Stroboscopic-Perturbation [541]. Strong [601]. Structure [713, 311, 354, 540, 571, 267, 80, 343, 398, 31, 617]. Studies [483, 286, 308, 237]. Study [456, 584, 34, 91, 105]. Studying [14]. Subgroups [632, 435]. Sufficient [596]. Sum [312, 379]. Summation [328, 606]. Sums [459, 191, 444]. Superconducting [241]. Superconductivity [348, 126, 650]. Superconductor [123]. Superfluid [352, 517, 518, 347]. Superfluidity [348]. Superselection [211]. Supplementary [620]. Support [153]. Surface [610, 649, 193, 261, 427]. Susceptibility [69, 242, 337]. Symbols [184]. Symmetric [162, 626, 442, 51, 184, 258, 145]. Symmetries [435, 565, 209]. Symmetry [615, 311, 381, 294, 671, 51, 141, 399, 376, 43]. Symmetry-Adapted [51, 141]. Synthetic [67]. System [382, 264, 215, 10, 11, 656, 527, 139, 391, 623, 464, 384, 390, 602, 285, 125, 169, 585, 73, 235, 266]. Systematic [603]. Systems [71, 463, 622, 352, 517, 518, 188, 189, 190, 411, 106, 483, 404, 314, 57, 442, 456, 535, 74, 196, 231, 232, 268, 428, 259, 52, 14, 168, 354, 412, 506, 559, 230, 423, 53, 100, 346].

Tables [19]. Tamburino [437]. TCP [45]. Technique [218, 13]. Techniques [284]. Temperature [192]. Temperatures [73, 266]. Tensor [726, 546, 238, 528, 158]. Tensors [727, 294, 655, 224, 708]. Terms [352, 596, 548, 687, 734, 735]. Test [525]. th [603]. th-Order [603]. Their [662, 129, 632]. Theorem [30, 250, 615, 25, 716, 549, 58, 413, 732, 422, 574, 165, 657, 87, 220, 383, 543]. Theorems [495, 68, 243, 489]. Theoretical [112, 532, 496, 688]. Theories [203, 556, 102, 671, 359, 532, 675]. Theory [30, 71, 658, 55, 77, 491, 324, 325, 634, 608, 345, 172, 298, 183, 484, 163, 188, 189, 190, 411, 381, 334, 284, 301, 379, 371, 402, 215, 216, 314, 309, 626, 400, 260, 200, 614, 619, 303, 431, 701, 342, 519, 520, 70, 157, 465, 393, 410, 244, 349, 511, 672, 453, 454, 535, 113, 722, 479, 268, 509, 604, 408, 593, 286, 308, 227, 116, 206, 540, 571, 199, 617, 513, 412, 492, 228, 305, 117, 159, 160, 650, 360, 386, 498, 563, 36, 565, 230, 276, 642, 271, 279, 558, 114, 144, 27, 50]. Theory [580, 175, 664, 154, 448, 68, 243, 31, 81, 67, 724, 460, 350, 601, 133, 122, 323, 723]. Thermal [701, 231, 346]. Thermally [625]. Thermodynamic [86, 266]. Thermodynamical [455]. Thermodynamics [340, 336, 533, 341]. Thermomagnetic [61]. Thermostatics [25]. Thin [427, 323]. Three [343, 543, 285, 258, 181, 403, 108]. Three-Body [285, 258]. Three-Dimensional [543, 108]. Three-Particle [343, 403]. Threefold [311]. Time [711, 357, 112, 219, 378, 97, 272]. Time- [112]. Time-Dependent [357]. Time-Energy [378]. Top [145]. Topological [662, 39]. Total [433, 689, 331]. Traces [186, 262]. Trajectories [639, 497]. Transformation [727]. Transformations [88, 204]. Transient [239, 171]. Transition [123, 3, 527, 118, 387, 241, 353]. Transitions [356]. Transmission [122]. Transparent [611]. Transport [696, 495, 163, 215, 216, 314, 102, 701, 722, 560, 513, 733, 589, 26, 131, 121]. Trapped [639]. Traveling [218]. Traveling-Wave [218]. Treating [541]. Treatment [112, 161]. Triality [671]. Triangular [547, 630]. Tridiagonal [149]. Truncated [400]. Truncation [26, 131]. Tschebyscheff [121]. Tube [9, 130]. Turning [666]. Twisted [130]. Two [20, 730, 494, 162, 173, 536, 384, 733, 589, 358, 715, 261, 478, 339, 562, 542, 544, 398, 643, 18, 724, 83, 122, 719, 720, 257, 107]. Two-Center [162, 544]. Two-Component [20]. Two-Dimensional [358, 261, 724, 719, 720, 107]. Two-Particle [384, 478]. Two-Point [398]. Two-Space [257]. Two-Wire [122]. Type [546, 245, 671, 179, 511, 4].

Uncertainty [248]. Undulatory [439]. Unified [329, 388, 157, 114]. Uniform [576, 737]. Unimodular [377, 481]. Uniqueness [401, 79, 495, 665, 339, 618]. Unitary [377, 451, 671, 481, 492, 461, 607, 707, 736, 480, 43]. Unstable [446]. Unti [437]. Unusual [166, 214]. Upper [635, 636]. Use [183, 69, 242, 378]. Used [81]. Useful [723]. Using [510, 569, 13]. Utilizing [168].

V [308, 412, 237]. Vacuum [400]. Validity [489]. Value [324, 325, 696, 438, 279, 81]. Valued [654]. Values [400]. Vanish [274]. Vanished [400]. Vapor [519, 520, 349, 350]. Vapor-Liquid [520, 350]. Variable [431]. Variables [640, 536, 502, 591, 562, 67]. Variation [702]. Variational [329, 388, 287, 352, 728, 296, 187, 722, 14, 4]. Varying [428]. Vector [634, 183, 135, 460]. Velocity [21]. Vertex [198, 523]. Very [417]. via [248]. Vibrating [14]. Vibrational [193, 540, 571, 194]. Vibrations [291]. Virial [288, 613]. Viscosity [216, 528]. Visited [467]. Volume [69, 242, 640, 68, 243].

Waals [519, 520, 349, 350]. Walk [467]. Walks [440, 644]. Wall [267]. Wannier [606]. Wave [583, 54, 661, 32, 365, 135, 210, 541, 116, 206, 126, 691, 35, 49, 160, 207, 218, 597, 595, 276, 670, 285, 306, 460]. Wave-Function [116]. Wavefunction [327]. Wavefunctions [712, 462, 553]. Waveguide [663]. Waves [610, 697, 682, 575, 663, 573, 710, 693, 255, 257, 119, 333, 740]. Way [311]. Weak [634, 359, 665, 648]. Weakly [265, 664]. Wedge [58]. Weights [619]. Were [717]. which [274]. whose [400]. Width [479, 480]. Wightman [152, 596, 273, 475]. Wigner [615, 110, 413]. Wire [122]. WKB [694].

X [182, 488]. X-Ray [182, 488].

Yeh [740]. Yevick [586]. Yukawa [32, 556].

Zero [40, 397, 592, 619, 616, 223, 687, 735, 266]. Zero-Mass [616]. Zero-Point [40, 397]. Zeros [716, 425].

References

Zumino:1960:GPP

[1] Bruno Zumino. Gauge properties of propagators in quantum electrodynamics. Journal of Mathematical Physics, 1(1):1–7, January 1960. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

     2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i1/p1_s1.

Toupin:1960:DCC

[2] R. A. Toupin and R. S. Rivlin. Dimensional changes in crystals caused by dislocations. Journal of Mathematical Physics, 1(1):8–15, January 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

     2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i1/p8_s1.

Erdelyi:1960:ASD

[3] A. Erdélyi. Asymptotic solutions of differential equations with transition points or singularities. Journal of Mathematical Physics, 1(1):16–26, January 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

     tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
     i1/p16_s1.

Ohmura:1960:MPH

[4] Takashi Ohmura. Minimum property in the Hulthén-type variational methods. Journal of Mathematical Physics, 1(1):27–34, January 1960.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i1/p27_s1.

Ohmura:1960:SEE

[5] Takashi Ohmura. Stationary expression for effective range. Journal of Mathematical Physics, 1(1):35–40, January 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v1/i1/p35_s1.

Klein:1960:MRP

[6] Abraham Klein. Mandelstam representation for potential scattering. Journal of Mathematical Physics, 1(1):41–47, January 1960. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v1/i1/p41_s1.

Gelfand:1960:IFS

[7] I. M. Gel’fand and A. M. Yaglom. Integration in functional spaces and its applications in quantum physics. Journal of Mathematical Physics, 1(1):48–69, January 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v1/i1/p48_s1.

Mazur:1960:PCE

[8] Peter Mazur and Elliott Montroll. Poincaré cycles, ergodicity, and irreversibility in assemblies of coupled harmonic oscillators. Journal of Mathematical Physics, 1(1):70–84, January 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i1/p70_s1.

Grossmann:1960:DET

[9] A. Grossmann. Description of the extended tube. Journal of Mathematical Physics, 1(2):85–86, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v1/i2/p85_s1.

Garcia-Colin:1960:PDFa

[10] Leopoldo S. Garcia-Colin. Pair distribution function of a hard sphere Bose system calculated by the pseudo-potential method. Journal of Mathematical Physics, 1(2):87–96, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i2/p87_s1.

Garcia-Colin:1960:PDFb

[11] Leopoldo S. Garcia-Colin and Jean Peretti. On the pair distribution function of a hard sphere Bose system. Journal of Mathematical Physics, 1 (2):97–106, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v1/i2/p97_s1.

VonRoos:1960:FSL

[12] Oldwig Von Roos. Formal solution of Liouville’s equation. Journal of Mathematical Physics, 1(2):107–111, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i2/p107_s1.

VonRoos:1960:SCB

[13] Oldwig Von Roos. Solution of the collisionless Boltzmann equation using a diagram technique. Journal of Mathematical Physics, 1(2):112– 120, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i2/p112_s1.

Luttinger:1960:VMS

[14] J. M. Luttinger and R. B. Thomas. Variational method for studying the motion of classical vibrating systems. Journal of Mathematical Physics, 1 (2):121–126, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v1/i2/p121_s1.

Pipkin:1960:ECD

[15] A. C. Pipkin and R. S. Rivlin. Electrical conduction in deformed isotropic materials. Journal of Mathematical Physics, 1(2):127–130, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i2/p127_s1.

Aaron:1960:CBE

[16] Ronald Aaron and Abraham Klein. Convergence of the Born expansion. Journal of Mathematical Physics, 1(2):131–138, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v1/i2/p131_s1.

Green:1960:RDP

[17] T. A. Green and O. E. Lanford. Rigorous derivation of the phase shift formula for the Hilbert space scattering operator of a single particle. Journal of Mathematical Physics, 1(2):139–148, February 1960. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v1/i2/p139_s1.

Tarski:1960:AFO

[18] Jan Tarski. Analyticity of the fourth order scattering amplitude with two complex invariants. Journal of Mathematical Physics, 1(2):149– 163, February 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i2/p149_s1.

Zak:1960:MOC

[19] J. Zak. Method to obtain the character tables of nonsymmorphic space groups. Journal of Mathematical Physics, 1(3):165–171, March 1960. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i3/p165_s1.

Bergmann:1960:PRT

[20] Otto Bergmann. Problem of Rainich for two-component spinors. Journal of Mathematical Physics, 1(3):172–177, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i3/p172_s1.

Backus:1960:LPOa

[21] George Backus. Linearized plasma oscillations in arbitrary electron velocity distributions. Journal of Mathematical Physics, 1(3):178–191, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i3/p178_s1.

Meeron:1960:NEI

[22] Emmanuel Meeron. Nodal expansions. III. Exact integral equations for particle correlation functions. Journal of Mathematical Physics, 1(3):192– 201, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i3/p192_s1.

Sherman:1960:CAI

[23] S. Sherman. Combinatorial aspects of the Ising model for ferromagnetism. I. A conjecture of Feynman on paths and graphs. Journal of Mathematical Physics, 1(3):202–217, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v1/i3/p202_s1. See addendum [470].

Prosperi:1960:ECQ

[24] G. M. Prosperi and A. Scotti. Ergodicity conditions in quantum mechanics. Journal of Mathematical Physics, 1(3):218–221, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v1/i3/p218_s1.

Bernstein:1960:PCL

[25] B. Bernstein. Proof of Carathéodory’s local theorem and its global application to thermostatics. Journal of Mathematical Physics, 1(3):222–224, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i3/p222_s1.

Wilf:1960:NIT

[26] Herbert S. Wilf. Numerical integration of the transport equation with no angular truncation. Journal of Mathematical Physics, 1(3):225–230, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/

i3/p225_s1. See errata [131].

Streater:1960:DCQa

[27] R. F. Streater. Double commutator in quantum field theory. Journal of Mathematical Physics, 1(3):231–233, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i3/p231_s1.

Lewis:1960:SRC

[28] Daniel C. Lewis and Olga Taussky. Some remarks concerning the real and imaginary parts of the characteristic roots of a finite matrix. Journal of Mathematical Physics, 1(3):234–236, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i3/p234_s1.

Lomont:1960:DDP

[29] J. S. Lomont. Decomposition of direct products of representations of the inhomogeneous Lorentz group. Journal of Mathematical Physics, 1(3): 237–243, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i3/p237_s1.

Albertoni:1960:NTC

[30] S. Albertoni, P. Bocchieri, and A. Loinger. New theorem in the classical ensemble theory. Journal of Mathematical Physics, 1(3):244–248, March 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i3/p244_s1.

Symanzik:1960:MPS

[31] K. Symanzik. On the many-particle structure of Green’s functions in quantum field theory. Journal of Mathematical Physics, 1(4):249–273, April 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i4/p249_s1.

Fivel:1960:APP

[32] Daniel I. Fivel and Abraham Klein. On the analytic properties of partial wave amplitudes in Yukawa potential scattering. Journal of Mathematical Physics, 1(4):274–279, April 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v1/i4/p274_s1.

Bellman:1960:IIM

[33] Richard Bellman, Robert Kalaba, and G. Milton Wing. Invariant imbedding and mathematical physics. I. Particle processes. Journal of Mathematical Physics, 1(4):280–308, April 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i4/p280_s1.

Rubin:1960:SDS

[34] Robert J. Rubin. Statistical dynamics of simple cubic lattices. model for the study of Brownian motion. Journal of Mathematical Physics, 1(4): 309–318, April 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i4/p309_s1. See errata [91].

Newton:1960:APRa

[35] Roger G. Newton. Analytic properties of radial wave functions. Journal of Mathematical Physics, 1(4):319–347, April 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i4/p319_s1.

Prigogine:1960:GTA

[36] I. Prigogine and F. Henin. On the general theory of the approach to equilibrium. I. Interacting normal modes. Journal of Mathematical Physics, 1(5):349–371, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v1/i5/p349_s1.

Teitler:1960:LER

[37] S. Teitler and R. F. Wallis. Liouville equation and the resolvent formalism. Journal of Mathematical Physics, 1(5):372–377, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v1/i5/p372_s1.

Siegel:1960:DOA

[38] Armand Siegel. Differential-operator approximations to the linear Boltzmann equation. Journal of Mathematical Physics, 1(5):378–390, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i5/p378_s1.

Green:1960:TDM

[39] Melville S. Green. Topological derivation of the Mayer density series for the pressure of an imperfect gas. Journal of Mathematical Physics, 1(5): 391–394, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i5/p391_s1.

Coldwell-Horsfall:1960:ZPE

[40] Rosemary A. Coldwell-Horsfall and Alexei A. Maradudin. Zero-point energy of an electron lattice. Journal of Mathematical Physics, 1(5):395–404, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/

i5/p395_s1. See erratum [397].

Sturrock:1960:GLE

[41] P. A. Sturrock. Generalization of the Lagrange expansion with application to physical problems. Journal of Mathematical Physics, 1(5):405–408, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i5/p405_s1.

Wigner:1960:NFA

[42] Eugene P. Wigner. Normal form of antiunitary operators. Journal of Mathematical Physics, 1(5):409–413, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i5/p409_s1.

Wigner:1960:PDB

[43] Eugene P. Wigner. Phenomenological distinction between unitary and antiunitary symmetry operators. Journal of Mathematical Physics, 1(5): 414–416, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i5/p414_s1.

Fano:1960:RRC

[44] U. Fano. Real representations of coordinate rotations. Journal of Mathematical Physics, 1(5):417–423, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i5/p417_s1.

Grossmann:1960:ACT

[45] A. Grossmann. Algebraic characterization of the TCP operation. Journal of Mathematical Physics, 1(5):424–428, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i5/p424_s1.

Cutkosky:1960:SDF

[46] R. E. Cutkosky. Singularities and discontinuities of Feynman amplitudes. Journal of Mathematical Physics, 1(5):429–433, May 1960. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v1/i5/p429_s1.

Arnowitt:1960:CCR

[47] R. Arnowitt, S. Deser, and C. W. Misner. Consistency of the canonical reduction of General Relativity. Journal of Mathematical Physics, 1(5): 434–439, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i5/p434_s1.

Finkelstein:1960:SEP

[48] R. Finkelstein. Spacetime of the elementary particles. Journal of Mathematical Physics, 1(5):440–451, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i5/p440_s1.

Newton:1960:APRb

[49] Roger G. Newton. Analytic properties of radial wave functions. Journal of Mathematical Physics, 1(5):452, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i5/p452_s2.

Streater:1960:DCQb

[50] R. F. Streater. Double commutator in quantum field theory. Journal of Mathematical Physics, 1(5):452, May 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i5/p452_s1.

McIntosh:1960:SAF

[51] Harold V. McIntosh. Symmetry-adapted functions belonging to the symmetric groups. Journal of Mathematical Physics, 1(6):453–460, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i6/p453_s1.

Lowdin:1960:CIO

[52] P.-O. Löwdin, R. Pauncz, and J. de Heer. On the calculation of the inverse of the overlap matrix in cyclic systems. Journal of Mathematical Physics, 1(6):461–467, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v1/i6/p461_s1.

Segal:1960:QNS

[53] I. E. Segal. Quantization of nonlinear systems. Journal of Mathematical Physics, 1(6):468–488, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v1/i6/p468_s1.

DaSilveira:1960:KWE

[54] Adel Da Silveira. Kemmer wave equation in Riemann space. Journal of Mathematical Physics, 1(6):489–491, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i6/p489_s1.

Araki:1960:HFC

[55] H. Araki. Hamiltonian formalism and the canonical commutation relations in quantum field theory. Journal of Mathematical Physics, 1(6):492–504, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i6/p492_s1.

Vosko:1960:ASB

[56] S. H. Vosko. Approximate solutions of the Bethe–Salpeter equation. Journal of Mathematical Physics, 1(6):505–515, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i6/p505_s1.

Girardeau:1960:RBS

[57] M. Girardeau. Relationship between systems of impenetrable bosons and fermions in one dimension. Journal of Mathematical Physics, 1(6): 516–523, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i6/p516_s1.

Epstein:1960:GEW

[58] H. Epstein. Generalization of the “edge-of-the-wedge” theorem. Journal of Mathematical Physics, 1(6):524–531, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i6/p524_s1.

Acharya:1960:FDR

[59] R. Acharya and E. C. G. Sudarshan. “front” description in relativistic quantum mechanics. Journal of Mathematical Physics, 1(6):532–536, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/
i6/p532_s1.

Buchdahl:1960:NCS

[60] H. A. Buchdahl. On the nonexistence of a class of static Einstein spaces asymptotic at infinity to a space of constant curvature. Journal of Mathematical Physics, 1(6):537–541, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v1/i6/p537_s1.

Pipkin:1960:GTE

[61] A. C. Pipkin and R. S. Rivlin. Galvanomagnetic and thermomagnetic effects in isotropic materials. Journal of Mathematical Physics, 1(6): 542–546, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i6/p542_s1.

Ehrman:1960:COP

[62] J. R. Ehrman, L. D. Fosdick, and D. C. Handscomb. Computation of order parameters in an Ising lattice by the Monte Carlo method. Journal of Mathematical Physics, 1(6):547–558, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v1/i6/p547_s1.

Backus:1960:LPOb

[63] George E. Backus. Linearized plasma oscillations in arbitrary electron distributions. Journal of Mathematical Physics, 1(6):559, June 1960. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v1/i6/p559_s1.

Saletan:1961:CLG

[64] Eugene J. Saletan. Contraction of Lie groups. Journal of Mathematical Physics, 2(1):1–21, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i1/p1_s1. See erratum [150].

Kursunoglu:1961:COA

[65] Behram Kursunoğlu. Complex orthogonal and antiorthogonal representation of Lorentz group. Journal of Mathematical Physics, 2(1):22–32, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i1/p22_s1.

Cook:1961:APB

[66] J. M. Cook. Asymptotic properties of a boson field with given source. Journal of Mathematical Physics, 2(1):33–45, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i1/p33_s1.

Tani:1961:SAC

[67] Smio Tani. Synthetic approach and canonical variables in a nonlocal field theory. Journal of Mathematical Physics, 2(1):46–51, January 1961. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i1/p46_s1.

Sykes:1961:SCT

[68] M. F. Sykes. Some counting theorems in the theory of the Ising model and the excluded volume problem. Journal of Mathematical Physics, 2(1): 52–62, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i1/p52_s1. See erratum [243].

Domb:1961:USE

[69] C. Domb and M. F. Sykes. Use of series expansions for the Ising model susceptibility and excluded volume problem. Journal of Mathematical Physics, 2(1):63–67, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i1/p63_s1. See erratum [242].

Henin:1961:GTA

[70] F. Henin, P. Résibois, and F. Andrews. On the general theory of the approach to equilibrium. II. Interacting particles. Journal of Mathematical Physics, 2(1):68–90, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i1/p68_s1.

Andrews:1961:GTA

[71] Frank C. Andrews. On the general theory of the approach to equilibrium. III. Inhomogeneous systems. Journal of Mathematical Physics, 2(1):91– 95, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i1/p91_s1.

Gioumousis:1961:MCIa

[72] George Gioumousis and C. F. Curtiss. Molecular collisions. II. Diatomic molecules. Journal of Mathematical Physics, 2(1):96–104, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i1/p96_s1.

Wu:1961:SNP

[73] Tai Tsun Wu. Some nonequilibrium properties of a Bose system of hard spheres at extremely low temperatures. Journal of Mathematical Physics, 2(1):105–123, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i1/p105_s1.

Kraichnan:1961:DNS

[74] Robert H. Kraichnan. Dynamics of nonlinear stochastic systems. Journal of Mathematical Physics, 2(1):124–148, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i1/p124_s1. See erratum [196].

Katz:1961:MBP

[75] Sylvan Katz. On a moving boundary problem. Journal of Mathematical Physics, 2(1):149–150, January 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i1/p149_s1.

DeWitt:1961:QFI

[76] Bryce S. DeWitt. Quantization of fields with infinite-dimensional invariance groups. Journal of Mathematical Physics, 2(2):151–162, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i2/p151_s1.

Araki:1961:GRF

[77] Huzihiro Araki. Generalized retarded functions and analytic function in momentum space in quantum field theory. Journal of Mathematical Physics, 2(2):163–177, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v2/i2/p163_s1.

Wichmann:1961:IRN

[78] Eyvind H. Wichmann and Ching-Hung Woo. Integral representation for the nonrelativistic Coulomb Green’s function. Journal of Mathematical Physics, 2(2):178–180, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v2/i2/p178_s1.

Barut:1961:APM

[79] A. O. Barut and K. H. Ruei. Analytical properties of S matrix and uniqueness of the scattering potential. Journal of Mathematical Physics, 2(2): 181–187, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i2/p181_s1.

Newton:1961:SMC

[80] Roger G. Newton. Structure of the many-channel S matrix. Journal of Mathematical Physics, 2(2):188–197, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i2/p188_s1.

Tani:1961:PPV

[81] S. Tani. Products of principal value singularities used in the formal theory of scattering. Journal of Mathematical Physics, 2(2):198–201, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i2/p198_s1.

Schell:1961:CFD

[82] J. F. Schell. Classification of four-dimensional Riemannian spaces. Journal of Mathematical Physics, 2(2):202–206, February 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i2/p202_s1.

Wigner:1961:GLP

[83] E. P. Wigner. Geometry of light paths between two material bodies. Journal of Mathematical Physics, 2(2):207–211, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i2/p207_s1.

Kibble:1961:LIG

[84] T. W. B. Kibble. Lorentz invariance and the gravitational field. Journal of Mathematical Physics, 2(2):212–221, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i2/p212_s1.

Lewis:1961:SES

[85] Robert M. Lewis. Solution of the equations of statistical mechanics. Journal of Mathematical Physics, 2(2):222–231, February 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i2/p222_s1.

Baym:1961:DTG

[86] Gordon Baym and N. David Mermin. Determination of thermodynamic Green’s functions. Journal of Mathematical Physics, 2(2):232–234, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i2/p232_s1.

Percival:1961:APQ

[87] Ian C. Percival. Almost periodicity and the quantal H theorem. Journal of Mathematical Physics, 2(2):235–239, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v2/i2/p235_s1. See addendum [220].

Bremermann:1961:ACM

[88] H. J. Bremermann and L. Durand. On analytic continuation, multiplication, and Fourier transformations of Schwartz distributions. Journal of Mathematical Physics, 2(2):240–258, February 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i2/p240_s1.

Conlan:1961:CI

[89] James Conlan, J. B. Diaz, and W. E. Parr. On the capacity of the icosahedron. Journal of Mathematical Physics, 2(2):259–261, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i2/p259_s1.

Keller:1961:LBI

[90] Joseph B. Keller. Lower bounds and isoperimetric inequalities for eigenvalues of the Schrödinger equation. Journal of Mathematical Physics, 2 (2):262–266, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i2/p262_s1.

Rubin:1961:ESD

[91] Robert J. Rubin. Errata: “Statistical Dynamics of Simple Cubic Lattices. Model for the Study of Brownian Motion”. Journal of Mathematical Physics, 2(2):266, February 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i2/p266_s1. See [34].

Araki:1961:CSC

[92] Huzihiro Araki. On the connection of spin and commutation relations between different fields. Journal of Mathematical Physics, 2(3):267–270, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i3/p267_s1.

Schwartz:1961:FQL

[93] J. Schwartz. Free quantized Lorentzian fields. Journal of Mathematical Physics, 2(3):271–290, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i3/p271_s1.

Robinson:1961:NEF

[94] Ivor Robinson. Null electromagnetic fields. Journal of Mathematical Physics, 2(3):290–291, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i3/p290_s1.

Layzer:1961:FPEa

[95] Arthur J. Layzer. Free propagator expansion in the evaluation of the Lamb shift. I. Journal of Mathematical Physics, 2(3):292–307, March 1961.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i3/p292_s1.

Layzer:1961:FPEb

[96] Arthur J. Layzer. Free propagator expansion in the evaluation of the Lamb shift. II. Journal of Mathematical Physics, 2(3):308–323, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i3/p308_s1.

Newman:1961:SPE

[97] E. Newman. Some properties of empty space-time. Journal of Mathematical Physics, 2(3):324–327, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i3/p324_s1.

Goldberg:1961:SAI

[98] J. N. Goldberg and R. P. Kerr. Some applications of the infinitesimalholonomy group to the Petrov classification of Einstein spaces. Journal of Mathematical Physics, 2(3):327–332, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i3/p327_s1.

Kerr:1961:ESF

[99] R. P. Kerr and J. N. Goldberg. Einstein spaces with four-parameter holonomy groups. Journal of Mathematical Physics, 2(3):332–336, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i3/p332_s1.

Temkin:1961:ADP

[100] A. Temkin. Absolute definition of phase shift in the elastic scattering of a particle from compound systems. Journal of Mathematical Physics, 2(3):336–340, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i3/p336_s1.

Lieb:1961:OFS

[101] Elliott Lieb. Operator formalism in statistical mechanics. Journal of Mathematical Physics, 2(3):341–343, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i3/p341_s1.

Green:1961:TTF

[102] H. S. Green. Theories of transport in fluids. Journal of Mathematical Physics, 2(3):344–348, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i3/p344_s1.

Maradudin:1961:NPI

[103] Alexei A. Maradudin, George H. Weiss, and Donald W. Jepsen. Nonequilibrium processes in isotopically disordered crystals. Journal of Mathematical Physics, 2(3):349–369, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i3/p349_s1.

Zwanzig:1961:NPI

[104] Robert Zwanzig. Nonequilibrium processes in isotopically disordered crystals. dependence on degree of order. Journal of Mathematical Physics, 2 (3):370–372, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i3/p370_s1.

Rubin:1961:SDS

[105] Robert J. Rubin. Statistical dynamics of simple cubic lattices. model for the study of Brownian motion. II. Journal of Mathematical Physics, 2(3): 373–386, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i3/p373_s1.

Ford:1961:EEN

[106] Joseph Ford. Equipartition of energy for nonlinear systems. Journal of Mathematical Physics, 2(3):387–393, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i3/p387_s1.

Zitron:1961:HOAa

[107] Norman Zitron and Samuel N. Karp. Higher-order approximations in multiple scattering. I. Two-dimensional scalar case. Journal of Mathematical Physics, 2(3):394–402, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i3/p394_s1.

Zitron:1961:HOAb

[108] Norman Zitron and Samuel N. Karp. Higher-order approximations in multiple scattering. II. Three-dimensional scalar case. Journal of Mathematical Physics, 2(3):402–406, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i3/p402_s1.

Schwinger:1961:BMQ

[109] Julian Schwinger. Brownian motion of a quantum oscillator. Journal of Mathematical Physics, 2(3):407–432, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i3/p407_s1.

Biedenharn:1961:WCG

[110] L. C. Biedenharn. Wigner coefficients for the R_{4} group and some applications. Journal of Mathematical Physics, 2(3):433–441, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i3/p433_s1.

Carlson:1961:EDC

[111] B. C. Carlson. Ellipsoidal distributions of charge or mass. Journal of Mathematical Physics, 2(3):441–450, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i3/p441_s1.

Guth:1961:GTT

[112] E. Guth and E. Inönü. Group-theoretical treatment of time- and energydependent multiple scattering with application to the slowing-down of neutrons. Journal of Mathematical Physics, 2(3):451–458, March 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i3/p451_s1.

Knight:1961:SLQ

[113] James M. Knight. Strict localization in quantum field theory. Journal of Mathematical Physics, 2(4):459–471, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i4/p459_s1.

Sciama:1961:IES

[114] D. W. Sciama. On the interpretation of the Einstein–Schrödinger unified field theory. Journal of Mathematical Physics, 2(4):472–477, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p472_s1.

Mapleton:1961:CCG

[115] Robert A. Mapleton. Characteristic Coulomb Green’s function and its eigenfunction expansion. Journal of Mathematical Physics, 2(4):478–482, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i4/p478_s1.

Mapleton:1961:ASW

[116] Robert A. Mapleton. Application of a singular wave-function operator in scattering theory. Journal of Mathematical Physics, 2(4):482–490, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i4/p482_s1.

OMalley:1961:MER

[117] Thomas F. O’Malley, Larry Spruch, and Leonard Rosenberg. Modification of effective-range theory in the presence of a long-range (r^{−4}) potential. Journal of Mathematical Physics, 2(4):491–498, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p491_s1.

Kowalski:1961:TMN

[118] K. L. Kowalski and D. Feldman. Transition matrix for nucleon-nucleon scattering. Journal of Mathematical Physics, 2(4):499–511, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p499_s1.

Waterman:1961:MSW

[119] P. C. Waterman and Rohn Truell. Multiple scattering of waves. Journal of Mathematical Physics, 2(4):512–537, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i4/p512_s1.

Zelazny:1961:ESC

[120] Roman Zelazny. Exact solution of a critical problem for a slab. Journal of Mathematical Physics, 2(4):538–542, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i4/p538_s1.

Yabushita:1961:TPA

[121] Shin Yabushita. Tschebyscheff polynomial approximation method of the neutron-transport equation. Journal of Mathematical Physics, 2(4):543– 549, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i4/p543_s1.

Wu:1961:TDA

[122] Tai Tsun Wu. Theory of the dipole antenna and the two-wire transmission line. Journal of Mathematical Physics, 2(4):550–574, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p550_s1.

Cohen:1961:BLB

[123] Hirsh Cohen and W. L. Miranker. Boundary-layer behavior in the superconductor transition problem. Journal of Mathematical Physics, 2(4): 575–583, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i4/p575_s1.

Langer:1961:FSD

[124] J. S. Langer. Frequency spectrum of a disordered one-dimensional lattice. Journal of Mathematical Physics, 2(4):584–591, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p584_s1.

VanKampen:1961:ECF

[125] N. G. Van Kampen. Exact calculation of the fluctuation spectrum for a nonlinear model system. Journal of Mathematical Physics, 2(4):592–601, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i4/p592_s1.

Mattis:1961:EWF

[126] Daniel Mattis and Elliott Lieb. Exact wave functions in superconductivity. Journal of Mathematical Physics, 2(4):602–609, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p602_s1.

Fisher:1961:SCS

[127] Michael E. Fisher and John W. Essam. Some cluster size and percolation problems. Journal of Mathematical Physics, 2(4):609–619, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p609_s1.

Fisher:1961:CPC

[128] Michael E. Fisher. Critical probabilities for cluster size and percolation problems. Journal of Mathematical Physics, 2(4):620–627, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p620_s1.

Frisch:1961:PSI

[129] Ivan T. Frisch and Wan H. Kim. Properties of 2-semi-isomorphic graphs and their applications: Active network analysis. Journal of Mathematical Physics, 2(4):627–635, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i4/p627_s1.

Pipkin:1961:ECS

[130] A. C. Pipkin and R. S. Rivlin. Electrical conduction in a stretched and twisted tube. Journal of Mathematical Physics, 2(4):636–638, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i4/p636_s1.

Wilf:1961:ENI

[131] Herbert S. Wilf. Errata: “Numerical Integration of the Transport Equation with No Angular Truncation”. Journal of Mathematical Physics, 2 (4):638, April 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i4/p638_s1. See [26].

Bros:1961:PAC

[132] J. Bros, A. Messiah, and R. Stora. A problem of analytic completion related to the Jost–Lehmann–Dyson formula. Journal of Mathematical Physics, 2(5):639–651, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i5/p639_s1.

Wellner:1961:RAC

[133] M. Wellner and R. B. Curtis. Role of the asymptotic condition in a Lagrangian field theory. Journal of Mathematical Physics, 2(5):651–655, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i5/p651_s1.

Eden:1961:ACL

[134] R. J. Eden, P. V. Landshoff, J. C. Polkinghorne, and J. C. Taylor. Acnodes and cusps on Landau curves. Journal of Mathematical Physics, 2(5):

656–663, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v2/i5/p656_s1.

Hjalmars:1961:WES

[135] Stig Hjalmars. Wave equations for scalar and vector particles in gravitational fields. Journal of Mathematical Physics, 2(5):663–666, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i5/p663_s1.

Newman:1961:NAEa

[136] E. T. Newman and L. A. Tamburino. New approach to Einstein’s empty space field equations. Journal of Mathematical Physics, 2(5):667–674, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i5/p667_s1.

Newman:1961:NAEb

[137] Ezra T. Newman. New approach to the Einstein and Maxwell-Einstein field equations. Journal of Mathematical Physics, 2(5):674–676, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i5/p674_s1.

Plebanski:1961:CC

[138] J. Plebanski and J. Ryten. Coordinate conditions. Journal of Mathematical Physics, 2(5):677–681, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i5/p677_s1.

Lenard:1961:ESM

[139] A. Lenard. Exact statistical mechanics of a one-dimensional system with Coulomb forces. Journal of Mathematical Physics, 2(5):682–693, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i5/p682_s1.

Carpenter:1961:GSH

[140] Chester P. Carpenter and Fritz W. Mezger. General spherical harmonics formulation of plasma Boltzmann equation. Journal of Mathematical Physics, 2(5):694–701, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i5/p694_s1.

Nesbet:1961:CSA

[141] R. K. Nesbet. Construction of symmetry-adapted functions in the manyparticle problem. Journal of Mathematical Physics, 2(5):701–709, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i5/p701_s1.

Grossmann:1961:SSAa

[142] Alex Grossmann and Tai Tsun Wu. Schrödinger scattering amplitude. I. Journal of Mathematical Physics, 2(5):710–713, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i5/p710_s1.

Grossmann:1961:SSAb

[143] Alex Grossmann. Schrödinger scattering amplitude. II. Journal of Mathematical Physics, 2(5):714–718, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i5/p714_s1.

Snowdon:1961:TMR

[144] S. C. Snowdon. Theory of multipole radiation. Journal of Mathematical Physics, 2(5):719–722, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i5/p719_s1.

Gioumousis:1961:MCIb

[145] George Gioumousis. Molecular collisions. III. Symmetric top molecules. Journal of Mathematical Physics, 2(5):723–727, May 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i5/p723_s1.

Hammersley:1961:CAB

[146] J. M. Hammersley. Comparison of atom and bond percolation processes. Journal of Mathematical Physics, 2(5):728–733, May 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i5/p728_s1.

Chasman:1961:EPM

[147] Richard R. Chasman. Eigenvalue problems in matrix mechanics. Journal of Mathematical Physics, 2(5):733–735, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i5/p733_s1.

Swalen:1961:RCF

[148] Jerome D. Swalen and Louis Pierce. Remarks on the continued fraction calculation of eigenvalues and eigenvectors. Journal of Mathematical Physics, 2(5):736–739, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i5/p736_s1.

Pierce:1961:CET

[149] Louis Pierce. Calculation of the eigenvalues of a tridiagonal Hermitian matrix. Journal of Mathematical Physics, 2(5):740–741, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i5/p740_s1.

Saletan:1961:ECL

[150] Eugene J. Saletan. Erratum: “Contraction of Lie Groups. Journal of Mathematical Physics, 2(5):742, May 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i5/p742_s1. See [64].

Lopuszanski:1961:CFC

[151] Jan Lopuszanski. A criterion for the free character of fields. Journal of Mathematical Physics, 2(6):743–747, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i6/p743_s1.

Fronsdal:1961:AWF

[152] Christian Fronsdal. Analyticity of Wightman functions. Journal of Mathematical Physics, 2(6):748–758, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i6/p748_s1.

DellAntonio:1961:SFS

[153] G. F. Dell’Antonio. Support of a field in p space. Journal of Mathematical Physics, 2(6):759–766, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i6/p759_s1.

Sudarshan:1961:NAR

[154] E. C. G. Sudarshan and K. Bardakci. Nature of the axioms of relativistic quantum field theory. I. Journal of Mathematical Physics, 2(6):767–771, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i6/p767_s1.

Jordan:1961:DMD

[155] Thomas F. Jordan and E. C. G. Sudarshan. Dynamical mappings of density operators in quantum mechanics. Journal of Mathematical Physics, 2(6):772–775, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i6/p772_s1.

Robinson:1961:NCM

[156] D. W. Robinson. A note on k-commutative matrices. Journal of Mathematical Physics, 2(6):776–777, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i6/p776_s1.

Israel:1961:NPU

[157] W. Israel and R. Trollope. New possibilities for a unified field theory. Journal of Mathematical Physics, 2(6):777–786, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v2/i6/p777_s1.

Taub:1961:ASE

[158] A. H. Taub. Approximate stress energy tensor for gravitational fields. Journal of Mathematical Physics, 2(6):787–793, June 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i6/p787_s1.

Odeh:1961:PLAa

[159] Farouk M. Odeh. Principles of limiting absorption and limiting amplitude in scattering theory. I. Schrödinger’s equation. Journal of Mathematical Physics, 2(6):794–800, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i6/p794_s1.

Odeh:1961:PLAb

[160] Farouk M. Odeh. Principles of limiting absorption and limiting amplitude in scattering theory. II. The wave equation in an inhomogeneous medium. Journal of Mathematical Physics, 2(6):800–802, June 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v2/i6/p800_s1.

Wakano:1961:FTS

[161] Masami Wakano. Foundations for a treatment of the scattering of light by the hydrodynamical and statistical atom model. Journal of Mathematical Physics, 2(6):803–824, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i6/p803_s1.

Fontana:1961:SEO

[162] Peter R. Fontana. Symmetric expansion of one- and two-center Coulomb potentials. Journal of Mathematical Physics, 2(6):825–828, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i6/p825_s1.

Dresner:1961:IOI

[163] Lawrence Dresner. Isoperimetric and other inequalities in the theory of neutron transport. Journal of Mathematical Physics, 2(6):829–847, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i6/p829_s1.

Broyles:1961:PDC

[164] A. A. Broyles. Probability distribution for classical fluids. Journal of Mathematical Physics, 2(6):848–857, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i6/p848_s1.

Lomont:1961:NAB

[165] J. S. Lomont. Note on the Albertoni–Bocchieri–Loinger theorem of classical statistical mechanics. Journal of Mathematical Physics, 2(6):858–860, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/
i6/p858_s1.

Rubin:1961:LUF

[166] Robert J. Rubin and Robert Zwanzig. A lattice with an unusual frequency spectrum. Journal of Mathematical Physics, 2(6):861–864, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i6/p861_s1.

Pipkin:1961:ECN

[167] A. C. Pipkin and R. S. Rivlin. Electrical conduction in a noncircular rod. Journal of Mathematical Physics, 2(6):865–868, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v2/i6/p865_s1.

Magiros:1961:MDP

[168] Demetrios G. Magiros. Method for defining principal modes of nonlinear systems utilizing infinite determinants. Journal of Mathematical Physics, 2(6):869–875, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v2/i6/p869_s1.

Wichmann:1961:NAA

[169] Eyvind H. Wichmann. Note on the algebraic aspect of the integration of a system of ordinary linear differential equations. Journal of Mathematical Physics, 2(6):876–880, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i6/p876_s1.

Struble:1961:GPS

[170] Raimond A. Struble and John E. Fletcher. General perturbational solution of the harmonically forced van der Pol equation. Journal of Mathematical Physics, 2(6):880–891, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v2/i6/p880_s1.

Wu:1961:TRD

[171] Tai Tsun Wu. Transient response of a dipole antenna. Journal of Mathematical Physics, 2(6):892–894, June 1961. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v2/i6/p892_s1.

Cook:1962:SPF

[172] L. F. Cook and Jan Tarski. Some properties of the five-point function in perturbation theory. Journal of Mathematical Physics, 3(1):1–30, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p1_s1.

Greenberg:1962:TFL

[173] O. W. Greenberg. Two folk lemmas on the expansion of the S matrix or the out field in normal ordered in fields. Journal of Mathematical Physics, 3(1):31–34, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i1/p31_s1.

Kerner:1962:HFA

[174] Edward H. Kerner. Hamiltonian formulation of action-at-a-distance in electrodynamics. Journal of Mathematical Physics, 3(1):35–42, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p35_s1.

Sturrock:1962:FTA

[175] P. A. Sturrock. Field-theory analogs of the Lagrange and Poincaré invariants. Journal of Mathematical Physics, 3(1):43–50, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p43_
s1.

Boulware:1962:SSE

[176] David G. Boulware. Spin and statistics with an electromagnetic field. Journal of Mathematical Physics, 3(1):50–58, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p50_s1.

See erratum [269].

Reiss:1962:ALL

[177] Howard R. Reiss. Absorption of light by light. Journal of Mathematical Physics, 3(1):59–67, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i1/p59_s1.

Ebel:1962:CRM

[178] M. E. Ebel. Causality and the R matrix. Journal of Mathematical Physics, 3(1):68–69, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i1/p68_s1.

Hale:1962:SPS

[179] Jack K. Hale and Arnold P. Stokes. Some physical solutions of Diractype equations. Journal of Mathematical Physics, 3(1):70–74, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p70_s1.

Newton:1962:CPP

[180] Roger G. Newton. Construction of potentials from the phase shifts at fixed energy. Journal of Mathematical Physics, 3(1):75–82, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p75_
s1.

Twersky:1962:MSA

[181] Victor Twersky. Multiple scattering by arbitrary configurations in three dimensions. Journal of Mathematical Physics, 3(1):83–91, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p83_s1.

Miller:1962:SAX

[182] Akeley Miller and Paul W. Schmidt. Small angle X-ray scattering from randomly oriented cylinders of arbitrary cross section. Journal of Mathematical Physics, 3(1):92–96, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i1/p92_s1.

Dedrick:1962:UDV

[183] K. G. Dedrick and Robert N. Wilson. Use of the displacement vector in electromagnetic theory. Journal of Mathematical Physics, 3(1):97–105, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i1/p97_s1.

Rose:1962:EES

[184] Morris E. Rose and C. N. Yang. Eigenvalues and eigenvectors of a symmetric matrix of 6 j symbols. Journal of Mathematical Physics, 3(1):106, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i1/p106_s1.

Gilbert:1962:IOM

[185] T. L. Gilbert. Inverse overlap matrix for periodic arrays of atoms. Journal of Mathematical Physics, 3(1):107–117, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p107_s1.

Ambler:1962:TPAa

[186] E. Ambler, J. C. Eisenstein, and J. F. Schooley. Traces of products of angular momentum matrices. Journal of Mathematical Physics, 3(1):118– 130, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i1/p118_s1.

Girardeau:1962:VMQ

[187] M. Girardeau. Variational method for the quantum statistics of interacting particles. Journal of Mathematical Physics, 3(1):131–139, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p131_s1. See erratum [296].

Dyson:1962:STEa

[188] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. I. Journal of Mathematical Physics, 3(1):140–156, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p140_s1.

Dyson:1962:STEb

[189] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. II. Journal of Mathematical Physics, 3(1):157–165, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p157_s1.

Dyson:1962:STEc

[190] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. III. Journal of Mathematical Physics, 3(1):166–175, January/ February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i1/p166_s1.

Rushbrooke:1962:CSI

[191] G. S. Rushbrooke and H. I. Scoins. Cluster sums for the Ising model. Journal of Mathematical Physics, 3(1):176–184, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p176_s1.

See errata [444].

Rushbrooke:1962:HTI

[192] G. S. Rushbrooke and J. Eve. High-temperature Ising partition function and related noncrossing polygons for the simple cubic lattice. Journal of Mathematical Physics, 3(1):185–189, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p185_s1.

Gazis:1962:SVM

[193] Denos C. Gazis and Richard F. Wallis. Surface vibrational modes in crystal lattices with complex interatomic interactions. Journal of Mathematical Physics, 3(1):190–199, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i1/p190_s1.

Rosenstock:1962:VMD

[194] Herbert B. Rosenstock and Richard E. McGill. Vibrational modes of disordered linear chains. Journal of Mathematical Physics, 3(1):200–202, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i1/p200_s1.

Blum:1962:ECO

[195] Lesser Blum. Evolution of a crystal in an oscillating field. Journal of Mathematical Physics, 3(1):203–204, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i1/p203_s1.

Kraichnan:1962:EDN

[196] Robert H. Kraichnan. Erratum: “Dynamics of Nonlinear Stochastic Systems”. Journal of Mathematical Physics, 3(1):205, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i1/p205_s1. See [74].

Finkelstein:1962:FQQ

[197] David Finkelstein, Josef M. Jauch, Samuel Schiminovich, and David Speiser. Foundations of quaternion quantum mechanics. Journal of Mathematical Physics, 3(2):207–220, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i2/p207_s1.

Brown:1962:APM

[198] W. S. Brown. Analyticity properties of the momentum-space vertex function. Journal of Mathematical Physics, 3(2):221–235, February 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p221_s1.

McKenna:1962:ARR

[199] James McKenna. Analytic regularization in relativistic quantum field theory. Journal of Mathematical Physics, 3(2):236–247, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p236_s1.

Haag:1962:PQF

[200] R. Haag and B. Schroer. Postulates of quantum field theory. Journal of Mathematical Physics, 3(2):248–256, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i2/p248_s1.

Streater:1962:APP

[201] R. F. Streater. Analytic properties of products of field operators. Journal of Mathematical Physics, 3(2):256–261, February 1962. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v3/i2/p256_s1.

Wu:1962:SPI

[202] Tai Tsun Wu. Some properties of impedance as a causal operator. Journal of Mathematical Physics, 3(2):262–271, February 1962. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v3/i2/p262_s1.

Frank:1962:CPE

[203] William M. Frank. Convergence of perturbation expansions in cutoff meson theories. Journal of Mathematical Physics, 3(2):272–280, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i2/p272_s1.

Halpern:1962:SEC

[204] Francis R. Halpern. Some examples of canonical transformations in quantum mechanics. Journal of Mathematical Physics, 3(2):281–286, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i2/p281_s1.

Bartram:1962:BEM

[205] Ralph Bartram and Larry Spruch. Bounds on elements of the S matrix for elastic scattering: One-dimensional scattering. Journal of Mathematical Physics, 3(2):287–296, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v3/i2/p287_s1.

Mapleton:1962:ASW

[206] Robert A. Mapleton. Application of the singular wave function operator in scattering theory. II. Journal of Mathematical Physics, 3(2):297– 304, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i2/p297_s1.

Pandres:1962:DAC

[207] Dave Pandres. Derivation of admissibility conditions for wave functions from general quantum-mechanical principles. Journal of Mathematical Physics, 3(2):305–308, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v3/i2/p305_s1.

Lumley:1962:AEL

[208] John L. Lumley. An approach to the Eulerian–Lagrangian problem. Journal of Mathematical Physics, 3(2):309–312, February 1962. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v3/i2/p309_s1.

Rosen:1962:SEM

[209] Gerald Rosen. Symmetries of the Einstein–Maxwell equations. Journal of Mathematical Physics, 3(2):313–318, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i2/p313_s1.

Johnson:1962:CCW

[210] W. R. Johnson and R. T. Deck. Continuum Coulomb wave function. Journal of Mathematical Physics, 3(2):319–324, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p319_s1.

Galindo:1962:SPP

[211] A. Galindo, A. Morales, and R. Nuñez-Lagos. Superselection principle and pure states of n-identical particles. Journal of Mathematical Physics, 3(2):324–328, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i2/p324_s1.

Prosperi:1962:ECQ

[212] G. M. Prosperi. Ergodicity conditions in quantum mechanics and Van Hove’s Hamiltonians. Journal of Mathematical Physics, 3(2):329–340, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i2/p329_s1.

Abraham:1962:CAF

[213] Phillip B. Abraham and George Weiss. On the calculation of analytic functions of cyclic matrices. Journal of Mathematical Physics, 3(2):340– 345, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i2/p340_s1.

daSilveira:1962:UEE

[214] Adel da Silveira. Unusual eigenvalues of the energy operator. Journal of Mathematical Physics, 3(2):346–349, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i2/p346_s1.

Fujita:1962:TTCa

[215] S. Fujita and R. Abe. Theory of transport coefficients. I. General theory and electrical conductivity of electron–phonon system. Journal of Mathematical Physics, 3(2):350–359, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i2/p350_s1.

Fujita:1962:TTCb

[216] S. Fujita. Theory of transport coefficients. II viscosity coefficients of quantum gases obeying the Boltzmann statistics. Journal of Mathematical Physics, 3(2):359–367, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v3/i2/p359_s1.

Pipkin:1962:NRC

[217] A. C. Pipkin and R. S. Rivlin. Non-rectilinear current flow in a straight conductor. Journal of Mathematical Physics, 3(2):368–371, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p368_s1.

Parzen:1962:TWT

[218] P. Parzen and G. Nomicos. Traveling-wave technique of cyclotron radiation. Journal of Mathematical Physics, 3(2):372–381, February 1962.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p372_s1.

Hori:1962:LTB

[219] Jun ichi Hori. Long-time behavior of a random lattice. Journal of Mathematical Physics, 3(2):382–386, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i2/p382_s1.

Percival:1962:AAP

[220] Ian C. Percival. Addendum: “Almost Periodicity and the Quantal H Theorem”. Journal of Mathematical Physics, 3(2):386, February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i2/p386_s1.

See [87].

Reiss:1962:CPE

[221] Howard R. Reiss. A convergent perturbation expansion in first-quantized electrodynamics. Journal of Mathematical Physics, 3(3):387–395, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i3/p387_s1.

Bowman:1962:GDCa

[222] John Judson Bowman and Joseph David Harris. Green’s distributions and the Cauchy problem for the iterated Klein–Gordon operator. Journal of Mathematical Physics, 3(3):396–404, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i3/p396_s1.

Lomont:1962:SRI

[223] J. S. Lomont and H. E. Moses. Simple realizations of the infinitesimal generators of the proper orthochronous inhomogeneous Lorentz group for mass zero. Journal of Mathematical Physics, 3(3):405–408, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p405_s1.

Rose:1962:PIA

[224] M. E. Rose. Properties of the irreducible angular momentum tensors. Journal of Mathematical Physics, 3(3):409–413, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p409_s1.

Jordan:1962:QMSa

[225] Thomas F. Jordan. The quantum-mechanical scattering problem. Journal of Mathematical Physics, 3(3):414–428, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i3/p414_s1.

Jordan:1962:QMSb

[226] Thomas F. Jordan. The quantum-mechanical scattering problem. II. Multi-channel scattering. Journal of Mathematical Physics, 3(3):429–439, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i3/p429_s1.

Lynch:1962:THE

[227] P. J. Lynch and B. C. Carlson. Theory of high-energy potential scattering. Journal of Mathematical Physics, 3(3):440–450, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p440_s1.

Murray:1962:PTL

[228] F. J. Murray. Perturbation theory and Lie algebras. Journal of Mathematical Physics, 3(3):451–468, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i3/p451_s1.

Bazley:1962:PEE

[229] Norman W. Bazley and David W. Fox. A procedure for estimating eigenvalues. Journal of Mathematical Physics, 3(3):469–471, March 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p469_s1.

Pytte:1962:NPT

[230] Agnar Pytte. A note on perturbation theory in nearly periodic systems. Journal of Mathematical Physics, 3(3):472–474, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p472_s1.

Kraichnan:1962:SMMa

[231] Robert H. Kraichnan. Stochastic models for many-body systems. I. Infinite systems in thermal equilibrium. Journal of Mathematical Physics, 3(3):475–495, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i3/p475_s1.

Kraichnan:1962:SMMb

[232] Robert H. Kraichnan. Stochastic models for many-body systems. II. Finite systems and statistical nonequilibrium. Journal of Mathematical Physics, 3(3):496–521, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i3/p496_s1.

Muhlschlegel:1962:AEB

[233] B. Mühlschlegel. Asymptotic expansion of the Bardeen–Cooper–Schrieffer partition function by means of the functional method. Journal of Mathematical Physics, 3(3):522–530, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i3/p522_s1.

Foldy:1962:AFS

[234] Leslie L. Foldy. Antisymmetric functions and Slater determinants. Journal of Mathematical Physics, 3(3):531–539, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i3/p531_s1.

Yanagawa:1962:ESF

[235] S. Yanagawa. Excitation spectrum of a fermion system. Journal of Mathematical Physics, 3(3):540–548, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i3/p540_s1.

Frisch:1962:LMS

[236] H. L. Frisch. Locally Maxwellian solutions for a “positronium” plasma. Journal of Mathematical Physics, 3(3):549, March 1962. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v3/i3/p549_s1.

Shuler:1962:SNR

[237] Kurt E. Shuler, George H. Weiss, and Knud Andersen. Studies in nonequilibrium rate processes. V. The relaxation of moments derived from a master equation. Journal of Mathematical Physics, 3(3):550–556, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p550_s1.

Judd:1962:DTO

[238] B. R. Judd. Double-tensor operators for configurations of equivalent electrons. Journal of Mathematical Physics, 3(3):557–563, March 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p557_s1.

Morgan:1962:TRD

[239] Samuel P. Morgan. Transient response of a dipole antenna. Journal of Mathematical Physics, 3(3):564–565, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i3/p564_s1.

Newman:1962:AGR

[240] Ezra Newman and Roger Penrose. An approach to gravitational radiation by a method of spin coefficients. Journal of Mathematical Physics, 3(3): 566–578, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i3/p566_s1. See errata [443].

Liniger:1962:STP

[241] Werner Liniger. The solution of a transition problem in a superconducting strip. Journal of Mathematical Physics, 3(3):578–586, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p578_s1.

Domb:1962:EIM

[242] C. Domb and M. F. Sykes. Erratum: “Ising Model Susceptibility [Use of Series Expansions for the Ising Model Susceptibility and Excluded Volume Problem]”. Journal of Mathematical Physics, 3(3):586, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i3/p586_s2.

See [69].

Sykes:1962:EIM

[243] M. F. Sykes. Erratum: “Ising Model and Excluded Volume Problem [Some Counting Theorems in the Theory of the Ising Model and the Excluded Volume Problem]”. Journal of Mathematical Physics, 3(3):586, March 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i3/p586_s1. See [68].

Jordan:1962:ROR

[244] Thomas F. Jordan and E. C. G. Sudarshan. Reduction of operator rings and the irreducibility axiom in quantum field theory. Journal of Mathematical Physics, 3(4):587–594, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p587_s1.

Fairlie:1962:SST

[245] D. B. Fairlie, P. V. Landshoff, J. Nuttall, and J. C. Polkinghorne. Singularities of the second type. Journal of Mathematical Physics, 3(4):594–602, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i4/p594_s1.

Pandres:1962:FIB

[246] Dave Pandres. On forces and interactions between fields. Journal of Mathematical Physics, 3(4):602–607, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p602_s1.

Utiyama:1962:RCG

[247] Ryoyu Utiyama and Bryce S. DeWitt. Renormalization of a classical gravitational field interacting with quantized matter fields. Journal of Mathematical Physics, 3(4):608–618, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p608_s1.

DeWitt:1962:DCU

[248] Bryce S. DeWitt. Definition of commutators via the uncertainty principle. Journal of Mathematical Physics, 3(4):619–624, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p619_s1.

DeWitt:1962:QFIa

[249] Bryce S. DeWitt. Quantization of fields with infinite-dimensional invariance groups. II. Anticommuting fields. Journal of Mathematical Physics, 3(4):625–636, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i4/p625_s1.

Arnowitt:1962:SST

[250] R. Arnowitt and S. Deser. The spin-statistics theorem. Journal of Mathematical Physics, 3(4):637–649, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p637_s1.

Kinoshita:1962:MSF

[251] Toichiro Kinoshita. Mass singularities of Feynman amplitudes. Journal of Mathematical Physics, 3(4):650–677, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i4/p650_s1.

Parrent:1962:MCF

[252] George B. Parrent, Robert A. Shore, and Thomas J. Skinner. On the mutual coherence function in an inhomogeneous medium. Journal of Mathematical Physics, 3(4):678–684, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p678_s1.

Grossmann:1962:SSA

[253] Alex Grossmann and Tai Tsun Wu. Schrödinger scattering amplitude. III. Journal of Mathematical Physics, 3(4):684–689, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p684_s1.

Meetz:1962:SSI

[254] K. Meetz. Spectrum of the scattering integral operator in the physical energy sheet. Journal of Mathematical Physics, 3(4):690–699, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p690_s1.

Twersky:1962:SWRa

[255] Victor Twersky. On scattering of waves by random distributions. I. Freespace scatterer formalism. Journal of Mathematical Physics, 3(4):700–715, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i4/p700_s1.

Twersky:1962:GCS

[256] Victor Twersky. On a general class of scattering problems. Journal of Mathematical Physics, 3(4):716–723, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p716_s1.

Twersky:1962:SWRb

[257] Victor Twersky. On scattering of waves by random distributions. II. Twospace scatterer formalism. Journal of Mathematical Physics, 3(4):724–734, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i4/p724_s1.

Smith:1962:SRT

[258] Felix T. Smith. A symmetric representation for three-body problems. I. Motion in a plane. Journal of Mathematical Physics, 3(4):735–748, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i4/p735_s1.

Lieb:1962:OEL

[259] Elliott Lieb and Daniel Mattis. Ordering energy levels of interacting spin systems. Journal of Mathematical Physics, 3(4):749–751, April 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p749_s1.

Gunson:1962:PCD

[260] J. Gunson. Proof of a conjecture by Dyson in the statistical theory of energy levels. Journal of Mathematical Physics, 3(4):752–753, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p752_s1.

Mullins:1962:PTD

[261] William W. Mullins. Proof that the two-dimensional shape of minimum surface free energy is convex. Journal of Mathematical Physics, 3(4): 754–759, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i4/p754_s1.

Ambler:1962:TPAb

[262] E. Ambler, J. C. Eisenstein, and J. F. Schooley. Traces of products of angular momentum matrices. II. Spherical basis. Journal of Mathematical Physics, 3(4):760–771, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i4/p760_s1.

Weiss:1962:BHF

[263] G. H. Weiss and A. A. Maradudin. The Baker–Hausdorff formula and a problem in crystal physics. Journal of Mathematical Physics, 3(4):

771–777, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i4/p771_s1.

Edwards:1962:ESM

[264] S. F. Edwards and A. Lenard. Exact statistical mechanics of a onedimensional system with Coulomb forces. II. The method of functional integration. Journal of Mathematical Physics, 3(4):778–792, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p778_s1.

Meijer:1962:DBS

[265] Paul H. E. Meijer, Tomoyasu Tanaka, and Jeremiah Barry. Dynamic behavior of a set of weakly coupled Ising spins. Journal of Mathematical Physics, 3(4):793–797, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i4/p793_s1.

Yang:1962:TFF

[266] Chen-Ping Yang. Thermodynamic functions of a Fermi system of dilute hard spheres at temperatures near absolute zero. Journal of Mathematical Physics, 3(4):797–800, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i4/p797_s1.

Muller:1962:SBW

[267] M. W. Muller and S. Dawson. Structure of a Bloch Wall. Journal of Mathematical Physics, 3(4):800–805, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i4/p800_s1.

Kruskal:1962:ATH

[268] Martin Kruskal. Asymptotic theory of Hamiltonian and other systems with all solutions nearly periodic. Journal of Mathematical Physics, 3(4): 806–828, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i4/p806_s1.

Boulware:1962:ESS

[269] David G. Boulware. Erratum: “Spin and Statistics with an Electromagnetic Field”. Journal of Mathematical Physics, 3(4):829, April 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i4/p829_s1.

See [176].

Schweber:1962:FQ

[270] S. S. Schweber. On Feynman quantization. Journal of Mathematical Physics, 3(5):831–842, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i5/p831_s1.

Sachs:1962:CFM

[271] Mendel Sachs and Solomon L. Schwebel. On covariant formulations of the Maxwell–Lorentz theory of electromagnetism. Journal of Mathematical Physics, 3(5):843–848, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i5/p843_s1.

Jordan:1962:DMD

[272] Thomas F. Jordan, Mark A. Pinsky, and E. C. G. Sudarshan. Dynamical mappings of density operators in quantum mechanics. II. Time dependent mappings. Journal of Mathematical Physics, 3(5):848–852, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p848_s1.

Manoharan:1962:PDH

[273] A. C. Manoharan. The primitive domains of holomorphy for the 4 and 5 point Wightman functions. Journal of Mathematical Physics, 3(5): 853–859, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i5/p853_s1.

Greenberg:1962:HFW

[274] O. W. Greenberg. Heisenberg fields which vanish on domains of momentum space. Journal of Mathematical Physics, 3(5):859–866, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p859_s1.

Newton:1962:NMP

[275] Roger G. Newton. Nonrelativistic S-matrix poles for complex angular momenta. Journal of Mathematical Physics, 3(5):867–882, May 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p867_s1.

See erratum [490].

Roos:1962:STD

[276] B. W. Roos and W. C. Sangren. Spectral theory of Dirac’s radial relativistic wave equation. Journal of Mathematical Physics, 3(5):882–890, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i5/p882_s1.

Newman:1962:BAF

[277] Ezra T. Newman and Theodore W. J. Unti. Behavior of asymptotically flat empty spaces. Journal of Mathematical Physics, 3(5):891–901, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p891_s1.

Newman:1962:ESM

[278] Ezra T. Newman and Louis A. Tamburino. Empty space metrics containing hypersurface orthogonal geodesic rays. Journal of Mathematical Physics, 3(5):902–907, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i5/p902_s1.

Sachs:1962:CIV

[279] R. K. Sachs. On the characteristic initial value problem in gravitational theory. Journal of Mathematical Physics, 3(5):908–914, May 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p908_s1.

Shale:1962:NSB

[280] David Shale. A note on the scattering of boson fields. Journal of Mathematical Physics, 3(5):915–921, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i5/p915_s1.

Froissart:1962:CAM

[281] Marcel Froissart. Complex angular momenta in potential scattering. Journal of Mathematical Physics, 3(5):922–927, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i5/p922_s1.

Percus:1962:EEA

[282] J. K. Percus and A. Rotenberg. Exact eigenfunctions of angular momentum by rotational projection. Journal of Mathematical Physics, 3(5):

928–932, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i5/p928_s1.

Kuroda:1962:PGL

[283] S. T. Kuroda. On a paper of Green and lanford. Journal of Mathematical Physics, 3(5):933–935, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i5/p933_s1.

Fowler:1962:IMS

[284] M. Fowler. Introduction to momentum space integration techniques in perturbation theory. Journal of Mathematical Physics, 3(5):936–945, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i5/p936_s1.

Slaggie:1962:APW

[285] E. Leo Slaggie and Eyvind H. Wichmann. Asymptotic properties of the wave function for a bound nonrelativistic three-body system. Journal of Mathematical Physics, 3(5):946–968, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i5/p946_s1.

Lowdin:1962:SPTa

[286] Per-Olov Löwdin. Studies in perturbation theory. IV. Solution of eigenvalue problem by projection operator formalism. Journal of Mathematical Physics, 3(5):969–982, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i5/p969_s1.

DeDominicis:1962:VFE

[287] Cyrano De Dominicis. Variational formulations of equilibrium statistical mechanics. Journal of Mathematical Physics, 3(5):983–1002, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p983_s1.

DeWitt:1962:APQ

[288] Hugh E. DeWitt. Analytic properties of the quantum corrections to the second virial coefficient. Journal of Mathematical Physics, 3(5):1003–1016, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i5/p1003_s1.

Swenson:1962:DGM

[289] Robert J. Swenson. Derivation of generalized master equations. Journal of Mathematical Physics, 3(5):1017–1022, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i5/p1017_s1.

Morrison:1962:NEL

[290] J. A. Morrison. On the number of electron levels in a one-dimensional random lattice. Journal of Mathematical Physics, 3(5):1023–1027, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p1023_s1.

Zabusky:1962:ESV

[291] Norman J. Zabusky. Exact solution for the vibrations of a nonlinear continuous model string. Journal of Mathematical Physics, 3(5):1028–1039, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i5/p1028_s1.

Wilson:1962:PCD

[292] Kenneth G. Wilson. Proof of a conjecture by Dyson. Journal of Mathematical Physics, 3(5):1040–1043, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i5/p1040_s1.

Abraham:1962:AFC

[293] Phillip B. Abraham and George H. Weiss. Analytic functions of continuant matrices. Journal of Mathematical Physics, 3(5):1044–1049, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/p1044_s1.

Grabner:1962:SRF

[294] Ludwig Grabner and John A. Swanson. Symmetry restrictions on field dependent tensors with application to galvanomagnetic effects. Journal of Mathematical Physics, 3(5):1050–1054, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i5/p1050_s1.

Zumino:1962:NFC

[295] Bruno Zumino. Normal forms of complex matrices. Journal of Mathematical Physics, 3(5):1055–1057, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i5/p1055_s1.

Girardeau:1962:EVM

[296] M. Girardeau. Erratum: “Variational Method for the Quantum Statistics of Interacting Particles”. Journal of Mathematical Physics, 3(5):1058, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i5/p1058_s1. See [187].

Gioumousis:1962:MCI

[297] George Gioumousis and C. F. Curtiss. Molecular collisions. IV. Nearly spherical rigid body approximation. Journal of Mathematical Physics, 3(6):1059–1072, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i6/p1059_s1.

DeWitt:1962:QFIb

[298] Bryce S. DeWitt. Quantization of fields with infinite-dimensional invariance groups. III. Generalized Schwinger–Feynman theory. Journal of Mathematical Physics, 3(6):1073–1093, November/December 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1073_s1.

Bialynicki-Birula:1962:GCQ

[299] I. Bialynicki-Birula. On the gauge covariance of quantum electrodynamics. Journal of Mathematical Physics, 3(6):1094–1098, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i6/p1094_s1.

Islam:1962:MMR

[300] Jamal N. Islam. Modified Mandelstam representation for heavy particles. Journal of Mathematical Physics, 3(6):1098–1106, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i6/p1098_s1.

Fried:1962:APT

[301] H. M. Fried. Axiomatic perturbation theory for retarded functions. Journal of Mathematical Physics, 3(6):1107–1115, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1107_s1.

Macfarlane:1962:RLG

[302] A. J. Macfarlane. On the restricted Lorentz group and groups homomorphically related to it. Journal of Mathematical Physics, 3(6):1116–1129, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1116_s1.

Halpern:1962:CSP

[303] Francis R. Halpern. A criterion for singularities in perturbation theory. Journal of Mathematical Physics, 3(6):1130–1134, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i6/p1130_s1.

Pais:1962:SD

[304] Abraham Pais. On spinors in n dimensions. Journal of Mathematical Physics, 3(6):1135–1139, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i6/p1135_s1.

Nakanishi:1962:IRP

[305] Noboru Nakanishi. Integral representations for production amplitudes in perturbation theory. Journal of Mathematical Physics, 3(6):1139–1146, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1139_s1.

Szasz:1962:AMB

[306] Levente Szász. Atomic many-body problem. II. The matrix components of the Hamiltonian with respect to correlated wave functions. Journal of Mathematical Physics, 3(6):1147–1160, November/December 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1147_s1.

Tani:1962:HCP

[307] S. Tani and D. A. Uhlenbrock. Hard core produced by orthogonality constraints. Journal of Mathematical Physics, 3(6):1161–1170, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i6/p1161_s1.

Lowdin:1962:SPTb

[308] Per-Olov Löwdin. Studies in perturbation theory. V. Some aspects on the exact self-consistent field theory. Journal of Mathematical Physics, 3(6):1171–1184, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v3/i6/p1171_s1.

Galindo:1962:FAM

[309] A. Galindo. On a functional averaging method in the classical ensemble theory. Journal of Mathematical Physics, 3(6):1185–1190, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i6/p1185_s1.

Dyson:1962:BMM

[310] Freeman J. Dyson. A Brownian-motion model for the eigenvalues of a random matrix. Journal of Mathematical Physics, 3(6):1191–1198, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i6/p1191_s1.

Dyson:1962:TWA

[311] Freeman J. Dyson. The threefold way. algebraic structure of symmetry groups and ensembles in quantum mechanics. Journal of Mathematical Physics, 3(6):1199–1215, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i6/p1199_s1.

DeWitt:1962:EQM

[312] Hugh E. DeWitt. Evaluation of the quantum-mechanical ring sum with Boltzmann statistics. Journal of Mathematical Physics, 3(6):1216–1228, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1216_s1.

Lewis:1962:QSB

[313] Robert M. Lewis. Quantum statistics and the Boltzmann equation. Journal of Mathematical Physics, 3(6):1229–1246, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1229_s1.

Fujita:1962:TTCc

[314] S. Fujita. Theory of transport coefficients. III. Quantum statistical systems. Journal of Mathematical Physics, 3(6):1246–1250, November/ December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i6/p1246_s1.

Fosdick:1962:NEP

[315] Lloyd D. Fosdick. Numerical estimation of the partition function in quantum statistics. Journal of Mathematical Physics, 3(6):1251–1264, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i6/p1251_s1.

Wu:1962:DRL

[316] Tai Tsun Wu. Dimers on rectangular lattices. Journal of Mathematical Physics, 3(6):1265–1266, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i6/p1265_s1.

Runge:1962:ASM

[317] Richard J. Runge. Approximate solutions of the modified Bloch equations for low magnetic fields. Journal of Mathematical Physics, 3(6):1267–1277, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1267_s1.

Zak:1962:SRI

[318] J. Zak. Selection rules for integrals of Bloch functions. Journal of Mathematical Physics, 3(6):1278–1279, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1278_s1.

Pasternack:1962:OPH

[319] S. Pasternack and R. M. Sternheimer. An orthogonality property of hydrogenlike radial functions. Journal of Mathematical Physics, 3(6):1280, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1280_s1.

Bowman:1962:GDCb

[320] John Judson Bowman and Joseph David Harris. Green’s distributions and the Cauchy problem for the multi-mass Klein–Gordon operator. Journal of Mathematical Physics, 3(6):1281–1290, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1281_s1.

Bowman:1962:GDA

[321] John Judson Bowman and Joseph David Harris. Green’s distributions associated with the operator [□^{m} − (−μ^{2})^{m}]^{l}. Journal of Mathematical Physics, 3(6):1291–1297, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v3/i6/p1291_s1.

Wu:1962:IAI

[322] Tai Tsun Wu. Input admittance of infinitely long dipole antennas driven from coaxial lines. Journal of Mathematical Physics, 3(6):1298–1301, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v3/i6/p1298_s1.

Wu:1962:TTC

[323] Tai Tsun Wu. Theory of the thin circular loop antenna. Journal of Mathematical Physics, 3(6):1301–1304, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1301_s1.

Bakshi:1963:EVFa

[324] Pradip M. Bakshi and Kalyana T. Mahanthappa. Expectation value formalism in quantum field theory. I. Journal of Mathematical Physics, 4(1): 1–11, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i1/p1_s1.

Bakshi:1963:EVFb

[325] Pradip M. Bakshi and Kalyana T. Mahanthappa. Expectation value formalism in quantum field theory. II. Journal of Mathematical Physics, 4(1):12–16, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i1/p12_s1.

Sucher:1963:RIS

[326] J. Sucher. Relativistic invariance and the square-root Klein–Gordon equation. Journal of Mathematical Physics, 4(1):17–23, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i1/p17_s1.

Sato:1963:NCB

[327] Iwao Sato. Normalization condition for the Bethe–Salpeter wavefunction and a formal solution to the Bethe–Salpeter equation. Journal of Mathematical Physics, 4(1):24–35, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i1/p24_s1.

Babbitt:1963:SPC

[328] Donald G. Babbitt. A summation procedure for certain Feynman integrals. Journal of Mathematical Physics, 4(1):36–41, January 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i1/p36_s1.

Corson:1963:UVFa

[329] Edward M. Corson. A unified variational formulation of classical and quantum dynamics. I. Journal of Mathematical Physics, 4(1):42–44, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i1/p42_s1.

Das:1963:CSF

[330] A. Das. Complex scalar field in general relativity. Journal of Mathematical Physics, 4(1):45–51, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i1/p45_s1.

Yang:1963:MDS

[331] C. N. Yang. Mathematical deductions from some rules concerning highenergy total cross sections. Journal of Mathematical Physics, 4(1):52– 53, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i1/p52_s1.

Levy:1963:LEE

[332] Bertram R. Levy and Joseph B. Keller. Low-energy expansion of scattering phase shifts for long-range potentials. Journal of Mathematical Physics, 4(1):54–64, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i1/p54_s1.

Yeh:1963:DWP

[333] C. Yeh. The diffraction of waves by a penetrable ribbon. Journal of Mathematical Physics, 4(1):65–71, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i1/p65_s1. See erratum [740].

Faddeyev:1963:IPQ

[334] L. D. Faddeyev and B. Seckler. The inverse problem in the quantum theory of scattering. Journal of Mathematical Physics, 4(1):72–104, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i1/p72_
s1.

Englman:1963:LSH

[335] Robert Englman and Peretz Levi. The line shape in a harmonic lattice. Journal of Mathematical Physics, 4(1):105–115, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i1/p105_s1.

Lebowitz:1963:STN

[336] J. L. Lebowitz and J. K. Percus. Statistical thermodynamics of nonuniform fluids. Journal of Mathematical Physics, 4(1):116–123, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i1/p116_s1.

Fisher:1963:PSI

[337] Michael E. Fisher. Perpendicular susceptibility of the Ising model. Journal of Mathematical Physics, 4(1):124–135, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i1/p124_s1.

Finkelstein:1963:QRC

[338] D. Finkelstein, J. M. Jauch, and D. Speiser. Quaternionic representations of Compact groups. Journal of Mathematical Physics, 4(1):136–140, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i1/p136_s1.

Odeh:1963:UES

[339] Farouk Odeh. Uniqueness and existence of the solution to the static London–Maxwell equations in two dimensions. Journal of Mathematical Physics, 4(1):141–146, January 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i1/p141_s1.

DeGroot:1963:DNT

[340] S. R. De Groot. On the development of nonequilibrium thermodynamics. Journal of Mathematical Physics, 4(2):147–153, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p147_s1.

Meixner:1963:TEN

[341] J. Meixner. Thermodynamics of electrical networks and the Onsager– Casimir reciprocal relations. Journal of Mathematical Physics, 4(2):154– 159, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i2/p154_s1.

Helfand:1963:TSE

[342] E. Helfand, R. J. Bearman, and V. S. Vaidhyanathan. Theory of the Soret effect in electrolytic solutions. Journal of Mathematical Physics, 4 (2):160–165, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i2/p160_s1.

Resibois:1963:STP

[343] P. Résibois. Structure of the three-particle scattering operator in classical gases. Journal of Mathematical Physics, 4(2):166–173, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p166_s1.

Kubo:1963:SLE

[344] Ryogo Kubo. Stochastic Liouville equations. Journal of Mathematical Physics, 4(2):174–183, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v4/i2/p174_s1.

Cohen:1963:KTD

[345] E. G. D. Cohen. On the kinetic theory of dense gases. Journal of Mathematical Physics, 4(2):183–189, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i2/p183_s1.

VanKampen:1963:TFN

[346] N. G. Van Kampen. Thermal fluctuations in nonlinear systems. Journal of Mathematical Physics, 4(2):190–194, February 1963. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i2/p190_s1.

Gross:1963:HSC

[347] Eugene P. Gross. Hydrodynamics of a superfluid condensate. Journal of Mathematical Physics, 4(2):195–207, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i2/p195_s1.

Martin:1963:MAS

[348] Paul C. Martin. A microscopic approach to superfluidity and superconductivity. Journal of Mathematical Physics, 4(2):208–215, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p208_s1.

Kac:1963:VWT

[349] M. Kac, G. E. Uhlenbeck, and P. C. Hemmer. On the van der Waals theory of the vapor–liquid equilibrium. I. Discussion of a one-dimensional model. Journal of Mathematical Physics, 4(2):216–228, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p216_s1.

Uhlenbeck:1963:VWT

[350] G. E. Uhlenbeck, P. C. Hemmer, and M. Kac. On the van der Waals theory of the vapor-liquid equilibrium. II. Discussion of the distribution functions. Journal of Mathematical Physics, 4(2):229–247, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p229_s1.

Lebowitz:1963:ABR

[351] J. L. Lebowitz and J. K. Percus. Asymptotic behavior of the radial distribution function. Journal of Mathematical Physics, 4(2):248–254, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i2/p248_s1.

DeDominicis:1963:VSM

[352] C. De Dominicis. Variational statistical mechanics in terms of “observables” for normal and superfluid systems. Journal of Mathematical Physics, 4(2):255–265, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v4/i2/p255_s1.

Verboven:1963:MES

[353] E. Verboven. The master equation with special transition probabilities. Journal of Mathematical Physics, 4(2):266–270, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i2/p266_s1.

Mazur:1963:AFS

[354] P. Mazur and J. Van Der Linden. Asymptotic form of the structure function for real systems. Journal of Mathematical Physics, 4(2):271– 277, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i2/p271_s1.

Fisher:1963:LSR

[355] Michael E. Fisher. Lattice statistics — a review and an exact isotherm for a plane lattice gas. Journal of Mathematical Physics, 4(2):278–286, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i2/p278_s1.

Kasteleyn:1963:DSP

[356] P. W. Kasteleyn. Dimer statistics and phase transitions. Journal of Mathematical Physics, 4(2):287–293, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i2/p287_s1.

Glauber:1963:TDS

[357] Roy J. Glauber. Time-dependent statistics of the Ising model. Journal of Mathematical Physics, 4(2):294–307, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i2/p294_s1.

Montroll:1963:CSM

[358] Elliott W. Montroll, Renfrey B. Potts, and John C. Ward. Correlations and spontaneous magnetization of the two-dimensional Ising model. Journal of Mathematical Physics, 4(2):308–322, February 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i2/p308_s1.

Haller:1963:SWD

[359] Kurt Haller. Scattering, weak decays, and final-state interactions in model theories in a dressed-particle picture. Journal of Mathematical Physics, 4(3):323–331, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i3/p323_s1.

Peres:1963:MMQ

[360] Asher Peres. Mechanical model for quantum field theory. Journal of Mathematical Physics, 4(3):332–333, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i3/p332_s1.

Betts:1963:ASN

[361] D. D. Betts, H. Schiff, and W. B. Strickfaden. Approximate solution of a nonlinear field equation. Journal of Mathematical Physics, 4(3):334– 338, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i3/p334_s1.

Robinson:1963:GMO

[362] Peter D. Robinson and Joseph O. Hirschfelder. Generalized momentum operators in quantum mechanics. Journal of Mathematical Physics, 4(3): 338–347, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i3/p338_s1.

Robinson:1963:NGM

[363] Peter D. Robinson. On the nonorthogonality of generalized momentum eigenfunctions in quantum mechanics. Journal of Mathematical Physics, 4(3):348–353, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i3/p348_s1.

Wybourne:1963:EIC

[364] B. G. Wybourne. Electrostatic interactions in complex electron configurations. Journal of Mathematical Physics, 4(3):354–356, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i3/p354_s1.

Hetherington:1963:PWA

[365] J. H. Hetherington. Partial wave analysis of the scattering of charged spinless particles. Journal of Mathematical Physics, 4(3):357–358, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i3/p357_s1.

Challifour:1963:RPB

[366] J. Challifour and R. J. Eden. Regge poles and branch cuts for potential scattering. Journal of Mathematical Physics, 4(3):359–371, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i3/p359_s1.

Belinfante:1963:PS

[367] Johan G. Belinfante and Burhan C. Unal. Potential scattering. Journal of Mathematical Physics, 4(3):372–387, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i3/p372_s1.

Dashen:1963:NAL

[368] Roger F. Dashen. New approach to low-energy potential scattering. Journal of Mathematical Physics, 4(3):388–391, March 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i3/p388_s1.

Rosen:1963:RCS

[369] Bernard Rosen. Relativistic Coulomb scattering. Journal of Mathematical Physics, 4(3):392–404, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i3/p392_s1.

Greenberg:1963:BKF

[370] Newton I. Greenberg. Binary kernel formulation of a Heisenberg model of ferromagnetism. Journal of Mathematical Physics, 4(3):405–409, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i3/p405_s1.

Frieman:1963:NMT

[371] Edward A. Frieman. On a new method in the theory of irreversible processes. Journal of Mathematical Physics, 4(3):410–418, March 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i3/p410_s1.

Yang:1963:SPR

[372] C. N. Yang. Some properties of the reduced density matrix. Journal of Mathematical Physics, 4(3):418–419, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i3/p418_s1.

Lomont:1963:EAI

[373] J. S. Lomont, H. E. Moses, and Kwang Yu. Equivalence and antiequivalence of irreducible sets of operators. I. Finite dimensional spaces. Journal of Mathematical Physics, 4(3):420–423, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i3/p420_s1.

Bragg:1963:SEF

[374] Lincoln E. Bragg and Bernard D. Coleman. On strain energy functions for isotropic elastic materials. Journal of Mathematical Physics, 4(3): 424–426, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i3/p424_s1.

Calogero:1963:NRE

[375] Francesco Calogero. Note on the Riccati equation. Journal of Mathematical Physics, 4(3):427–430, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i3/p427_s1.

Trammell:1963:SFC

[376] George T. Trammell. Symmetry functions of the cube. Journal of Mathematical Physics, 4(3):431–436, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i3/p431_s1.

Biedenharn:1963:RSL

[377] L. C. Biedenharn. On the representations of the semisimple Lie groups. I. The explicit construction of invariants for the unimodular unitary group in N dimensions. Journal of Mathematical Physics, 4(3):436–445, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i3/p436_s1.

Leibowitz:1963:MNT

[378] Martin A. Leibowitz. Moments of the neutron time-energy distribution and the use of random functionals. Journal of Mathematical Physics, 4(3): 446–450, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i3/p446_s1.

Fried:1963:APT

[379] H. M. Fried. Axiomatic perturbation theory. II. Partial sum formalism. Journal of Mathematical Physics, 4(4):451–464, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p451_s1.

Amar:1963:GPF

[380] V. Amar, M. Lucchini, and P. G. Sona. Group properties of free-field equations. Journal of Mathematical Physics, 4(4):465–468, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p465_s1.

Emch:1963:SQT

[381] Gérard Emch and Constantin Piron. Symmetry in quantum theory. Journal of Mathematical Physics, 4(4):469–473, April 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i4/p469_s1.

Aviles:1963:BBS

[382] Joseph B. Aviles. N-body Bose system with a finite number of states. I. Irreducible representations. Journal of Mathematical Physics, 4(4):473–483, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i4/p473_s1.

Robinson:1963:GTG

[383] I. Robinson and A. Schild. Generalization of a theorem by Goldberg and sachs. Journal of Mathematical Physics, 4(4):484–489, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p484_s1.

Macfarlane:1963:KRT

[384] Alan J. Macfarlane. Kinematics of the relativistic two-particle system. Journal of Mathematical Physics, 4(4):490–500, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p490_s1.

Goldberg:1963:DGM

[385] H. Goldberg and Y. Lehrer-Ilamed. Derivation of the Gell-Mann–Okubo mass formula. Journal of Mathematical Physics, 4(4):501–502, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p501_s1.

Polkinghorne:1963:HEBa

[386] J. C. Polkinghorne. High-energy behavior in perturbation theory. Journal of Mathematical Physics, 4(4):503–506, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i4/p503_s1.

Kowalski:1963:TMN

[387] K. L. Kowalski and D. Feldman. Transition matrix for nucleon-nucleon scattering. II. Journal of Mathematical Physics, 4(4):507–518, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p507_s1.

Corson:1963:UVFb

[388] Edward M. Corson. Unified variational formulation of classical and quantum dynamics. II. Journal of Mathematical Physics, 4(4):519–521, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i4/p519_s1.

Mano:1963:EFN

[389] Koichi Mano. Eigenfunction form of the nonrelativistic Coulomb Green’s function. Journal of Mathematical Physics, 4(4):522–529, April 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p522_s1.

Parry:1963:SAP

[390] W. E. Parry and R. E. Turner. Some analytic properties of Green functions and self-energy parts for a system of interacting bosons. Journal of Mathematical Physics, 4(4):530–533, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i4/p530_s1.

Lenard:1963:ESM

[391] Andrew Lenard. Exact statistical mechanics of a one-dimensional system with Coulomb forces. III. Statistics of the electric field. Journal of Mathematical Physics, 4(4):533–543, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i4/p533_s1.

Swenson:1963:GME

[392] Robert J. Swenson. Generalized master equation and t-matrix expansion. Journal of Mathematical Physics, 4(4):544–551, April 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p544_s1.

Jackson:1963:NCOa

[393] E. Atlee Jackson. Nonlinear coupled oscillators. I. Perturbation theory; ergodic problem. Journal of Mathematical Physics, 4(4):551–558, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i4/p551_s1.

Fowler:1963:LSC

[394] T. K. Fowler. Lyapunov’s stability criteria for plasmas. Journal of Mathematical Physics, 4(4):559–569, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i4/p559_s1.

Tarski:1963:PFC

[395] Jan Tarski. Partition function for certain simple Lie algebras. Journal of Mathematical Physics, 4(4):569–574, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i4/p569_s1.

Wei:1963:LAS

[396] James Wei and Edward Norman. Lie algebraic solution of linear differential equations. Journal of Mathematical Physics, 4(4):575–581, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i4/p575_s1.

Coldwell-Horsfall:1963:EZP

[397] Rosemary A. Coldwell-Horsfall and Alexei A. Maradudin. Erratum: “Zero-Point Energy of an Electron Lattice”. Journal of Mathematical Physics, 4(4):582, April 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i4/p582_s1. See [40].

Steinmann:1963:STP

[398] O. Steinmann. Structure of the two-point function. Journal of Mathematical Physics, 4(5):583–588, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i5/p583_s1.

Speiser:1963:PSG

[399] David R. Speiser and Jan Tarski. Possible schemes for global symmetry. Journal of Mathematical Physics, 4(5):588–612, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i5/p588_s1.

Greenberg:1963:QFT

[400] O. W. Greenberg and A. L. Licht. Quantum field-theory model whose truncated vacuum expectation values vanished beyond some order. Journal of Mathematical Physics, 4(5):613–614, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i5/p613_s1.

Arnowitt:1963:NUC

[401] R. Arnowitt and S. Deser. Note on uniqueness of canonical commutation relations. Journal of Mathematical Physics, 4(5):615–617, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i5/p615_s1.

Fuchs:1963:HEB

[402] N. H. Fuchs. High-energy behavior of a certain class of scattering amplitudes in perturbation theory. Journal of Mathematical Physics, 4(5): 617–621, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i5/p617_s1.

Whitten:1963:TPS

[403] R. C. Whitten. Three-particle scattering. I. Planar case. Journal of Mathematical Physics, 4(5):622–633, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i5/p622_s1.

Fredkin:1963:AFP

[404] Donald R. Fredkin. Asymptotic form of the proper self-energy function in many-particle systems. Journal of Mathematical Physics, 4(5):634–636, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i5/p634_s1.

Araki:1963:RCC

[405] H. Araki and E. J. Woods. Representations of the canonical commutation relations describing a nonrelativistic infinite free Bose gas. Journal of Mathematical Physics, 4(5):637–662, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i5/p637_s1.

Derrick:1963:RAO

[406] G. H. Derrick. Representation of annihilation operators in the space of periodic functions. Journal of Mathematical Physics, 4(5):663–666, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i5/p663_s1.

Schultz:1963:NOD

[407] T. D. Schultz. Note on the one-dimensional gas of impenetrable pointparticle bosons. Journal of Mathematical Physics, 4(5):666–671, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i5/p666_s1.

Lieb:1963:NMT

[408] Elliott Lieb. New method in the theory of imperfect gases and liquids. Journal of Mathematical Physics, 4(5):671–678, May 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i5/p671_s1.

Ninham:1963:AFC

[409] B. W. Ninham. Asymptotic form of the coefficients of some Ising-model series. Journal of Mathematical Physics, 4(5):679–685, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i5/p679_s1.

Jackson:1963:NCOb

[410] E. Atlee Jackson. Nonlinear coupled oscillators. II. Comparison of theory with computer solutions. Journal of Mathematical Physics, 4(5):686–700, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i5/p686_s1.

Dyson:1963:STE

[411] Freeman J. Dyson and Madan Lal Mehta. Statistical theory of the energy levels of complex systems. IV. Journal of Mathematical Physics, 4(5): 701–712, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v4/i5/p701_s1;   http://link.aip.org/link/?JMP/4/701/
1.

Mehta:1963:STE

[412] Madan Lal Mehta and Freeman J. Dyson. Statistical theory of the energy levels of complex systems. V. Journal of Mathematical Physics, 4(5): 713–719, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i5/p713_s1;   http://link.aip.org/link/?JMP/4/713/
1.

Ginibre:1963:WET

[413] Jean Ginibre. Wigner–Eckart theorem and simple Lie groups. Journal of Mathematical Physics, 4(5):720–726, May 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i5/p720_s1.

Fried:1963:PGR

[414] H. M. Fried and J. G. Taylor. Propagator ghosts and Regge poles. Journal of Mathematical Physics, 4(6):727–730, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i6/p727_s1.

Pagels:1963:DCG

[415] Heinz R. Pagels. Divergence conditions and the gravitational field equations. Journal of Mathematical Physics, 4(6):731–734, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p731_s1.

Manasse:1963:FNC

[416] F. K. Manasse and C. W. Misner. Fermi normal coordinates and some basic concepts in differential geometry. Journal of Mathematical Physics, 4(6):735–745, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i6/p735_s1.

Manasse:1963:DMS

[417] F. K. Manasse. Distortion in the metric of a small center of gravitational attraction due to its proximity to a very large mass. Journal of Mathematical Physics, 4(6):746–761, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i6/p746_s1.

Aizu:1963:PDQ

[418] Kêitsiro Aizu. Parameter differentiation of quantum-mechanical linear operators. Journal of Mathematical Physics, 4(6):762–775, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p762_s1.

Levy-Leblond:1963:GGN

[419] Jean-Marc Levy-Leblond. Galilei group and nonrelativistic quantum mechanics. Journal of Mathematical Physics, 4(6):776–788, June 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p776_s1.

Finkelstein:1963:PGQ

[420] D. Finkelstein, J. M. Jauch, S. Schiminovich, and D. Speiser. Principle of general Q covariance. Journal of Mathematical Physics, 4(6):788–796, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i6/p788_s1.

Syed:1963:RME

[421] A. Syed. Reduction of S-matrix elements. Journal of Mathematical Physics, 4(6):797–808, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i6/p797_s1.

Iwamoto:1963:TES

[422] Fumiaki Iwamoto. Theorem on energy shifts due to a potential in a large box. Journal of Mathematical Physics, 4(6):809–811, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p809_s1.

Ron:1963:GDF

[423] Amiram Ron. Generalized dielectric function for classical many-body systems. Journal of Mathematical Physics, 4(6):811–823, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p811_s1.

Taylor:1963:LLS

[424] Leonard S. Taylor. Lorentz–Lorenz and Sellmeier formulas in irregular gases. Journal of Mathematical Physics, 4(6):824–828, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i6/p824_s1.

Keller:1963:ZHF

[425] Joseph B. Keller, S. I. Rubinow, and Max Goldstein. Zeros of Hankel functions and poles of scattering amplitudes. Journal of Mathematical Physics, 4(6):829–832, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i6/p829_s1.

Smythe:1963:CSC

[426] W. R. Smythe. Charged spheroid in cylinder. Journal of Mathematical Physics, 4(6):833–837, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i6/p833_s1.

Serbin:1963:QIS

[427] Hyman Serbin. Quadratic invariants of surface deformations and the strain energy of thin elastic shells. Journal of Mathematical Physics, 4(6): 838–851, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i6/p838_s1.

Leibowitz:1963:SBL

[428] M. A. Leibowitz. Statistical behavior of linear systems with randomly varying parameters. Journal of Mathematical Physics, 4(6):852–858, June 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i6/p852_s1.

Fronsdal:1963:MIR

[429] C. Fronsdal, R. E. Norton, and K. T. Mahanthappa. More integral representations for scattering amplitudes with complex singularities. Journal of Mathematical Physics, 4(7):859–871, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i7/p859_s1.

Islam:1963:ACM

[430] Jamal N. Islam. Acnodes and cusps and the Mandelstam representation. Journal of Mathematical Physics, 4(7):872–878, July 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i7/p872_s1.

Halpern:1963:SFP

[431] Francis R. Halpern. Singularities of the four-point function in perturbation theory with variable masses. Journal of Mathematical Physics, 4(7):

879–884, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i7/p879_s1.

Wichmann:1963:DMA

[432] Eyvind H. Wichmann. Density matrices arising from incomplete measurements. Journal of Mathematical Physics, 4(7):884–896, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i7/p884_s1.

Granzow:1963:DTO

[433] Kenneth D. Granzow. N-dimensional total orbital angular-momentum operator. Journal of Mathematical Physics, 4(7):897–900, July 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i7/p897_s1.

Dirac:1963:RRS

[434] P. A. M. Dirac. A remarkable representation of the 3 + 2 de Sitter group. Journal of Mathematical Physics, 4(7):901–909, July 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i7/p901_s1.

Peaslee:1963:PCS

[435] D. C. Peaslee. Particle charge symmetries as R_{8} subgroups. Journal of Mathematical Physics, 4(7):910–914, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i7/p910_s1.

Newman:1963:ESG

[436] E. Newman, L. Tamburino, and T. Unti. Empty-space generalization of the Schwarzschild metric. Journal of Mathematical Physics, 4(7):915–923, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i7/p915_s1.

Misner:1963:FRN

[437] Charles W. Misner. The flatter regions of Newman, Unti, and Tamburino’s generalized Schwarzschild space. Journal of Mathematical Physics, 4(7): 924–937, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i7/p924_s1.

Lindquist:1963:IVP

[438] Richard W. Lindquist. Initial-value problem on Einstein–Rosen manifolds. Journal of Mathematical Physics, 4(7):938–950, July 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i7/p938_s1.

Lanczos:1963:URS

[439] C. Lanczos. Undulatory Riemannian spaces. Journal of Mathematical Physics, 4(7):951–959, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i7/p951_s1.

Kesten:1963:NSA

[440] Harry Kesten. On the number of self-avoiding walks. Journal of Mathematical Physics, 4(7):960–969, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i7/p960_s1.

Sasakawa:1963:NMS

[441] Tatuya Sasakawa. New method for solving eigenvalue problems. Journal of Mathematical Physics, 4(7):970–992, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i7/p970_s1.

Harker:1963:SCP

[442] Kenneth J. Harker. Solution of the Cauchy problem for Laplace’s equation in axially symmetric systems. Journal of Mathematical Physics, 4(7): 993–997, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i7/p993_s1.

Newman:1963:EAG

[443] E. Newman and R. Penrose. Errata: “An Approach to Gravitational Radiation by a Method of Spin Coefficients”. Journal of Mathematical Physics, 4(7):998, July 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i7/p998_s1. See [240].

Rushbrooke:1963:ECS

[444] G. S. Rushbrooke and H. I. Scoins. Errata: “Cluster Sums for the Ising Model”. Journal of Mathematical Physics, 4(7):998, July 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i7/p998_s2.

See [191].

Roos:1963:EAP

[445] B. W. Roos and W. C. Sangren. Expansions associated with a pair of singular first-order differential equations. Journal of Mathematical Physics, 4(8):999–1008, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i8/p999_s1.

Balescu:1963:KEU

[446] R. Balescu. Kinetic equation for an unstable plasma. Journal of Mathematical Physics, 4(8):1009–1019, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i8/p1009_s1.

Pagnamenta:1963:RMC

[447] Antonio Pagnamenta. Regularity of the T matrix in the case of Dirac potential scattering. Journal of Mathematical Physics, 4(8):1020–1028, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i8/p1020_s1.

Sudarshan:1963:NAR

[448] E. C. G. Sudarshan. The nature of the axioms of relativistic quantum field theory. II. Journal of Mathematical Physics, 4(8):1029–1036, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i8/p1029_s1.

Karr:1963:GCC

[449] Philip R. Karr. Generalized “cross-correlation” field quantities as solutions of field equations. Journal of Mathematical Physics, 4(8):1037–1039, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i8/p1037_s1.

Porter:1963:RMD

[450] C. E. Porter. Random matrix diagonalization-some numerical computations. Journal of Mathematical Physics, 4(8):1039–1042, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i8/p1039_s1.

Chan:1963:RUR

[451] Hong-Mo Chan. Real unitary representation of the many-channel S matrix for complex l and E. Journal of Mathematical Physics, 4(8):1042– 1047, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i8/p1042_s1.

Prosser:1963:RPS

[452] Reese T. Prosser. Relativistic potential scattering. Journal of Mathematical Physics, 4(8):1048–1054, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i8/p1048_s1.

Klauder:1963:CRTa

[453] John R. Klauder. Continuous-representation theory. I. Postulates of continuous-representation theory. Journal of Mathematical Physics, 4(8): 1055–1058, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i8/p1055_s1.

Klauder:1963:CRTb

[454] John R. Klauder. Continuous-representation theory. II. Generalized relation between quantum and classical dynamics. Journal of Mathematical Physics, 4(8):1058–1073, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v4/i8/p1058_s1.

Bragg:1963:TLC

[455] Lincoln E. Bragg and Bernard D. Coleman. A thermodynamical limitation on compressibility. Journal of Mathematical Physics, 4(8):1074–1077, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i8/p1074_s1.

Kac:1963:SSL

[456] M. Kac and E. Helfand. Study of several lattice systems with long-range forces. Journal of Mathematical Physics, 4(8):1078–1088, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i8/p1078_s1.

Jordan:1963:IRG

[457] T. F. Jordan, N. Mukunda, and S. V. Pepper. Irreducible representations of generalized oscillator operators. Journal of Mathematical Physics, 4 (8):1089–1095, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i8/p1089_s1.

Girardeau:1963:FMB

[458] M. Girardeau. Formulation of the many-body problem for composite particles. Journal of Mathematical Physics, 4(8):1096–1115, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i8/p1096_s1.

Abraham:1963:ECS

[459] Philip B. Abraham and George H. Weiss. Evaluation of certain sums arising in chemical physics. Journal of Mathematical Physics, 4(8):1115– 1118, August 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i8/p1115_s1.

Teitler:1963:AWV

[460] S. Teitler. Asymptotic wave vector and nonrelativistic perturbation theory. Journal of Mathematical Physics, 4(9):1119–1127, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1119_s1.

Moshinsky:1963:BIR

[461] Marcos Moshinsky. Bases for the irreducible representations of the unitary groups and some applications. Journal of Mathematical Physics, 4(9): 1128–1139, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i9/p1128_s1.

Sasaki:1963:SCA

[462] Fukashi Sasaki and Kimio Ohno. Spin-component analysis of singledeterminant wavefunctions. Journal of Mathematical Physics, 4(9):1140– 1147, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i9/p1140_s1.

Bazley:1963:LBE

[463] Norman W. Bazley and David W. Fox. Lower bounds for energy levels of molecular systems. Journal of Mathematical Physics, 4(9):1147– 1153, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i9/p1147_s1.

Luttinger:1963:ESM

[464] J. M. Luttinger. An exactly soluble model of a many-fermion system. Journal of Mathematical Physics, 4(9):1154–1162, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1154_s1.

Israel:1963:RKT

[465] Werner Israel. Relativistic kinetic theory of a simple gas. Journal of Mathematical Physics, 4(9):1163–1181, September 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1163_s1.

Ron:1963:KEE

[466] Amiram Ron. Kinetic equations for electrons and phonons. Journal of Mathematical Physics, 4(9):1182–1190, September 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1182_s1.

Vineyard:1963:NDS

[467] George H. Vineyard. The number of distinct sites visited in a random walk on a lattice. Journal of Mathematical Physics, 4(9):1191–1193, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i9/p1191_s1.

Metzner:1963:OPL

[468] A. W. K. Metzner. Observable properties of large relativistic masses. Journal of Mathematical Physics, 4(9):1194–1205, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1194_s1.

Islam:1963:MAE

[469] Mohammed Azizul Islam. Mathematical analysis on the effect of a prolate spheroidal core in a magnetic dipole field. Journal of Mathematical Physics, 4(9):1206–1212, September 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i9/p1206_s1.

Sherman:1963:ACA

[470] S. Sherman. Addendum: “Combinatorial Aspects of the Ising Model for Ferromagnetism. I. A Conjecture of Feynman on Paths and Graphs”.

Journal of Mathematical Physics, 4(9):1213–1214, September 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i9/p1213_s1.

See [23].

Chakrabarti:1963:CFC

[471] A. Chakrabarti. Canonical form of the covariant free-particle equations. Journal of Mathematical Physics, 4(10):1215–1222, October 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1215_
s1.

Chakrabarti:1963:RPO

[472] A. Chakrabarti. Relativistic position operator for free particles. Journal of Mathematical Physics, 4(10):1223–1228, October 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1223_s1.

Nakanishi:1963:ISE

[473] Noboru Nakanishi. Invariant solutions of the exact Bethe–Salpeter equation in the general-mass case. Journal of Mathematical Physics, 4(10): 1229–1235, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i10/p1229_s1.

Nakanishi:1963:RDD

[474] Noboru Nakanishi. Remarks on the double dispersion approach to the Bethe–Salpeter equation. Journal of Mathematical Physics, 4(10):1235– 1240, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i10/p1235_s1.

Tomozawa:1963:LCA

[475] Yukio Tomozawa. Local commutativity and the analytic continuation of the Wightman function. Journal of Mathematical Physics, 4(10):1240– 1252, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i10/p1240_s1.

Bessis:1963:CSP

[476] D. Bessis and F. Pham. Complex singularities in production amplitudes. Journal of Mathematical Physics, 4(10):1253–1257, October 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1253_
s1.

Tani:1963:RPH

[477] S. Tani. Regge poles in high-energy potential scattering. Journal of Mathematical Physics, 4(10):1258–1262, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i10/p1258_s1.

Nesbet:1963:CME

[478] R. K. Nesbet. Computation of matrix elements of integrable nuclear twoparticle interaction operators. Journal of Mathematical Physics, 4(10): 1262–1271, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i10/p1262_s1.

Krieger:1963:RWA

[479] T. J. Krieger and C. E. Porter. Reduced width amplitude distributions and random sign rules in R-matrix theory. Journal of Mathematical Physics, 4(10):1272–1279, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i10/p1272_s1.

Ullah:1963:RWA

[480] Nazakat Ullah. Reduced width amplitude distributions for the unitary ensemble. Journal of Mathematical Physics, 4(10):1279–1282, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1279_
s1.

Klein:1963:IOU

[481] Abraham Klein. Invariant operators of the unitary unimodular group in n dimensions. Journal of Mathematical Physics, 4(10):1283–1284, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i10/p1283_s1.

Bednarz:1963:EDB

[482] Roman J. Bednarz and Janusz R. Mika. Energy-dependent Boltzmann equation in plane geometry. Journal of Mathematical Physics, 4(10):1285– 1292, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i10/p1285_s1.

Ford:1963:CSE

[483] Joseph Ford and John Waters. Computer studies of energy sharing and ergodicity for nonlinear oscillator systems. Journal of Mathematical Physics, 4(10):1293–1306, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i10/p1293_s1.

Dimmock:1963:RTN

[484] John O. Dimmock. Representation theory for nonunitary groups. Journal of Mathematical Physics, 4(10):1307–1311, October 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1307_s1.

Penrose:1963:CFE

[485] Oliver Penrose. Convergence of fugacity expansions for fluids and lattice gases. Journal of Mathematical Physics, 4(10):1312–1320, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1312_
s1.

Burgoyne:1963:RCA

[486] P. N. Burgoyne. Remarks on the combinatorial approach to the Ising problem. Journal of Mathematical Physics, 4(10):1320–1326, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1320_
s1.

Rosen:1963:AEF

[487] Gerald Rosen. Approximate evaluation of Feynman functional integrals. Journal of Mathematical Physics, 4(10):1327–1333, October 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1327_
s1.

Rothwell:1963:NSA

[488] W. S. Rothwell. New small-angle X-ray scattering calculation. Journal of Mathematical Physics, 4(10):1334–1336, October 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i10/p1334_s1.

Wei:1963:NGV

[489] James Wei. Note on the global validity of the Baker–Hausdorff and Magnus theorems. Journal of Mathematical Physics, 4(10):1337–1341, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i10/p1337_s1.

Newton:1963:ENM

[490] Roger G. Newton. Errata: “Nonrelativistic S-Matrix Poles for Complex Angular Momenta”. Journal of Mathematical Physics, 4(10):1342, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i10/p1342_s1. See [275].

Araki:1963:LNA

[491] Huzihiro Araki. A lattice of von Neumann algebras associated with the quantum theory of a free Bose field. Journal of Mathematical Physics, 4(11):1343–1362, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i11/p1343_s1.

Moravek:1963:MCU

[492] Paul H. Moravek and David W. Joseph. Method for construction of unitary operators in quantum field theory. Journal of Mathematical Physics, 4(11):1363–1366, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i11/p1363_s1.

Andersen:1963:MLE

[493] Hans C. Andersen and Irwin Oppenheim. Modified Langevin equation for the description of Brownian motion. Journal of Mathematical Physics, 4(11):1367–1372, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i11/p1367_s1.

Chambers:1963:CTG

[494] Ll. G. Chambers. Comparison of two generalizations of Maxwell’s equations involving creation of charge. Journal of Mathematical Physics, 4 (11):1373–1375, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i11/p1373_s1.

Case:1963:EUT

[495] K. M. Case and P. F. Zweifel. Existence and uniqueness theorems for the neutron transport equation. Journal of Mathematical Physics, 4(11): 1376–1385, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i11/p1376_s1.

Nakanishi:1963:FPP

[496] Noboru Nakanishi. Fundamental properties of perturbation-theoretical integral representations. II. Journal of Mathematical Physics, 4(11):1385– 1392, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i11/p1385_s1.

Polkinghorne:1963:SRT

[497] J. C. Polkinghorne. Singularities of Regge trajectories and asymptotes to Landau curves. Journal of Mathematical Physics, 4(11):1393–1395, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i11/p1393_s1.

Polkinghorne:1963:HEBb

[498] J. C. Polkinghorne. High-energy behavior in perturbation theory. II. Journal of Mathematical Physics, 4(11):1396–1400, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v4/i11/p1396_s1.

Barut:1963:BSA

[499] A. O. Barut and J. Dilley. Behavior of the scattering amplitude for large angular momentum. Journal of Mathematical Physics, 4(11):1401– 1408, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i11/p1401_s1.

Sartori:1963:ABS

[500] L. Sartori. Asymptotic behavior of Schrödinger scattering amplitudes. Journal of Mathematical Physics, 4(11):1408–1414, November 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i11/p1408_
s1.

Dullemond:1963:CSE

[501] C. Dullemond. Coupled Schrödinger equations and statistical boundary conditions. Journal of Mathematical Physics, 4(11):1415–1432, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i11/p1415_s1.

Kibble:1963:CVI

[502] T. W. B. Kibble. Canonical variables for the interacting gravitational and Dirac fields. Journal of Mathematical Physics, 4(11):1433–1437, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i11/p1433_s1.

Wu:1963:CDB

[503] F. Y. Wu. Cluster development in an N-body problem. Journal of Mathematical Physics, 4(11):1438–1443, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v4/i11/p1438_s1.

Licht:1963:SL

[504] A. L. Licht. Strict localization. Journal of Mathematical Physics, 4(11): 1443–1447, November 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i11/p1443_s1.

Baird:1963:RSL

[505] G. E. Baird and L. C. Biedenharn. On the representations of the semisimple Lie groups. II. Journal of Mathematical Physics, 4(12):1449–1466, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i12/p1449_s1.

Newman:1963:CNF

[506] Ezra T. Newman and T. W. J. Unti. A class of null flat-space coordinate systems. Journal of Mathematical Physics, 4(12):1467–1469, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i12/p1467_s1.

Currie:1963:ICC

[507] Douglas G. Currie. Interaction contra classical relativistic Hamiltonian particle mechanics. Journal of Mathematical Physics, 4(12):1470–1488, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i12/p1470_s1.

Penrose:1963:RMF

[508] Oliver Penrose. The remainder in Mayer’s fugacity series. Journal of Mathematical Physics, 4(12):1488–1494, December 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v4/i12/p1488_s1.

Lebowitz:1963:IEI

[509] J. L. Lebowitz and J. K. Percus. Integral equations and inequalities in the theory of fluids. Journal of Mathematical Physics, 4(12):1495–1506, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i12/p1495_s1.

Chisholm:1963:SLI

[510] J. S. R. Chisholm. Solution of linear integral equations using Padé approximants. Journal of Mathematical Physics, 4(12):1506–1510, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i12/p1506_s1.

Kadison:1963:RTN

[511] Richard V. Kadison. Remarks on the type of von Neumann algebras of local observables in quantum field theory. Journal of Mathematical Physics, 4(12):1511–1516, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i12/p1511_s1.

Erma:1963:PAE

[512] Victor A. Erma. Perturbation approach to the electrostatic problem for irregularly shaped conductors. Journal of Mathematical Physics, 4(12): 1517–1526, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i12/p1517_s1.

McLennan:1963:TTC

[513] James A. McLennan and Robert J. Swenson. Theory of transport coefficients in low-density gases. Journal of Mathematical Physics, 4(12): 1527–1536, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v4/i12/p1527_s1.

Brysk:1963:DSF

[514] Henry Brysk. Determinantal solution of the Fredholm equation with Green’s function kernel. Journal of Mathematical Physics, 4(12):1536– 1538, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v4/i12/p1536_s1.

Nakanishi:1963:EMS

[515] Noboru Nakanishi. External-mass singularity. Journal of Mathematical Physics, 4(12):1539–1541, December 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i12/p1539_s1.

Araki:1964:NAL

[516] Huzihiro Araki. Von Neumann algebras of local observables for free scalar field. Journal of Mathematical Physics, 5(1):1–13, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p1_s1.

DeDominicis:1964:SEPa

[517] Cyrano De Dominicis and Paul C. Martin. Stationary entropy principle and renormalization in normal and superfluid systems. I. Algebraic formulation. Journal of Mathematical Physics, 5(1):14–30, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p14_s1.

DeDominicis:1964:SEPb

[518] Cyrano De Dominicis and Paul C. Martin. Stationary entropy principle and renormalization in normal and superfluid systems. II. Diagrammatic formulation. Journal of Mathematical Physics, 5(1):31–59, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p31_s1.

Hemmer:1964:VWTa

[519] P. C. Hemmer, M. Kac, and G. E. Uhlenbeck. On the van der Waals theory of the vapor–liquid equilibrium. III. Discussion of the critical region. Journal of Mathematical Physics, 5(1):60–74, January 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i1/p60_s1.

Hemmer:1964:VWTb

[520] P. C. Hemmer. On the van der Waals theory of the vapor-liquid equilibrium. IV. The pair correlation function and equation of state for longrange forces. Journal of Mathematical Physics, 5(1):75–84, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p75_s1.

Peterson:1964:GME

[521] Robert L. Peterson and Paul M. Quay. Generalized master equations for the density matrix. Journal of Mathematical Physics, 5(1):85–89, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p85_
s1.

Hurst:1964:APM

[522] C. A. Hurst. Applicability of the Pfaffian method to combinatorial problems on a lattice. Journal of Mathematical Physics, 5(1):90–100, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p90_
s1.

Fronsdal:1964:IRV

[523] C. Fronsdal and R. E. Norton. Integral representations for vertex functions. Journal of Mathematical Physics, 5(1):100–108, January 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p100_s1.

Bacry:1964:PPO

[524] H. Bacry. Position and polarization operators in relativistic and nonrelativistic mechanics. Journal of Mathematical Physics, 5(1):109–111, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i1/p109_s1.

Taub:1964:MTB

[525] A. H. Taub. Motion of test bodies in general relativity. Journal of Mathematical Physics, 5(1):112–119, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i1/p112_s1.

Dragt:1964:PCS

[526] A. J. Dragt and R. Karplus. Poles in coupled scattering amplitudes. Journal of Mathematical Physics, 5(1):120–126, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p120_s1.

Helfand:1964:APT

[527] Eugene Helfand. Approach to a phase transition in a one-dimensional system. Journal of Mathematical Physics, 5(1):127–132, January 1964.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p127_s1.

Schulz:1964:TAV

[528] William D. Schulz. Tensor artificial viscosity for numerical hydrodynamics. Journal of Mathematical Physics, 5(1):133–138, January 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p133_s1.

Struble:1964:MRD

[529] R. A. Struble and T. C. Harris. Motion of a relativistic damped oscillator. Journal of Mathematical Physics, 5(1):138–141, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p138_s1.

Soda:1964:ESI

[530] Toshio Soda. Excitation spectrum of an impurity in the hard-sphere Bose gas. Journal of Mathematical Physics, 5(1):142–146, January 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i1/p142_s1.

See errata [628].

Sooy:1964:ACS

[531] Walter R. Sooy. Approximate calculations of scattering phase shifts. Journal of Mathematical Physics, 5(1):147–154, January 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i1/p147_s1.

Joos:1964:GTM

[532] Hans Joos. Group-theoretical models of local-field theories. Journal of Mathematical Physics, 5(2):155–164, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i2/p155_s1.

Mandelbrot:1964:DST

[533] Benoit Mandelbrot. On the derivation of statistical thermodynamics from purely phenomenological principles. Journal of Mathematical Physics, 5 (2):164–171, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p164_s1.

Goldberg:1964:APE

[534] J. N. Goldberg and R. P. Kerr. Asymptotic properties of the electromagnetic field. Journal of Mathematical Physics, 5(2):172–176, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i2/p172_s1.

Klauder:1964:CRT

[535] John R. Klauder. Continuous-representation theory. III. On functional quantization of classical systems. Journal of Mathematical Physics, 5 (2):177–187, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p177_s1.

Gunther:1964:SAV

[536] Marian Günther. Separation of angular variables in the Bethe–Salpeter equation for two spin-1/2 particles. Journal of Mathematical Physics, 5 (2):188–192, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p188_s1.

Bjorken:1964:RBF

[537] James D. Bjorken. Regge behavior of forward elastic scattering amplitudes. Journal of Mathematical Physics, 5(2):192–198, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i2/p192_s1.

Morse:1964:DPC

[538] Burt Jules Morse. Diffraction by polygonal cylinders. Journal of Mathematical Physics, 5(2):199–214, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i2/p199_s1.

Bronk:1964:ASA

[539] Burt V. Bronk. Accuracy of the semicircle approximation for the density of eigenvalues of random matrices. Journal of Mathematical Physics, 5 (2):215–220, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p215_s1.

McCumber:1964:TVSa

[540] D. E. McCumber. Theory of vibrational structure in optical spectra of impurities in solids. I. Singlets. Journal of Mathematical Physics, 5(2):

221–230, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p221_s1.

Kruskal:1964:SPP

[541] Martin D. Kruskal and Norman J. Zabusky. Stroboscopic-perturbation procedure for treating a class of nonlinear wave equations. Journal of Mathematical Physics, 5(2):231–244, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i2/p231_s1.

Sack:1964:GLE

[542] R. A. Sack. Generalization of Laplace’s expansion to arbitrary powers and functions of the distance between two points. Journal of Mathematical Physics, 5(2):245–251, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i2/p245_s1.

Sack:1964:TDA

[543] R. A. Sack. Three-dimensional addition theorem for arbitrary functions involving expansions in spherical harmonics. Journal of Mathematical Physics, 5(2):252–259, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i2/p252_s1.

Sack:1964:TCE

[544] R. A. Sack. Two-center expansion for the powers of the distance between two points. Journal of Mathematical Physics, 5(2):260–268, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i2/p260_s1.

Segal:1964:EFC

[545] I. E. Segal. Explicit formal construction of nonlinear quantum fields. Journal of Mathematical Physics, 5(2):269–282, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i2/p269_s1.

Chiu:1964:ITE

[546] Ying-Nan Chiu. Irreducible tensor expansion of solid spherical harmonictype operators in quantum mechanics. Journal of Mathematical Physics, 5 (2):283–288, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i2/p283_s1.

Rosen:1964:SPT

[547] S. P. Rosen. Some properties of triangular representations of SU(3). Journal of Mathematical Physics, 5(2):289–293, February 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i2/p289_s1.

Lomont:1964:RILa

[548] J. S. Lomont and H. E. Moses. The representations of the inhomogeneous Lorentz group in terms of an angular momentum basis. Journal of Mathematical Physics, 5(2):294–298, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i2/p294_s1. See errata [734].

Cannon:1964:NIT

[549] John T. Cannon and Thomas F. Jordan. A no-interaction theorem in classical relativistic Hamiltonian particle dynamics. Journal of Mathematical Physics, 5(3):299–307, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i3/p299_s1.

Bevensee:1964:QES

[550] R. M. Bevensee. Quantum electrodynamics of the stimulated emission of radiation. I. Journal of Mathematical Physics, 5(3):308–324, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i3/p308_s1.

Rohrlich:1964:FDC

[551] F. Rohrlich. Functional differential calculus of operators. Journal of Mathematical Physics, 5(3):324–331, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i3/p324_s1.

Nelson:1964:FIS

[552] Edward Nelson. Feynman integrals and the Schrödinger equation. Journal of Mathematical Physics, 5(3):332–343, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i3/p332_s1.

Yeh:1964:ASC

[553] C. Yeh. An application of Sommerfeld’s complex-order wavefunctions to an antenna problem. Journal of Mathematical Physics, 5(3):344–350, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i3/p344_s1.

Durso:1964:RPP

[554] John W. Durso and Peter Signell. Regge poles and potentials with cores. Journal of Mathematical Physics, 5(3):350–354, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i3/p350_s1.

Garrido:1964:GAI

[555] L. M. Garrido. Generalized adiabatic invariance. Journal of Mathematical Physics, 5(3):355–362, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i3/p355_s1.

Frank:1964:CYT

[556] William M. Frank. Convergence of Yukawa theories with a finite number of interacting boson modes. Journal of Mathematical Physics, 5(3):363– 372, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i3/p363_s1.

Havas:1964:CBC

[557] Peter Havas. The connection between conservation laws and laws of motion in affine spaces. Journal of Mathematical Physics, 5(3):373–378, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i3/p373_s1.

Sasakawa:1964:DMP

[558] Tatuya Sasakawa. Determinantal method in perturbation theory. Journal of Mathematical Physics, 5(3):379–382, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i3/p379_s1.

Northcote:1964:ESE

[559] Robert S. Northcote and Renfrey B. Potts. Energy sharing and equilibrium for nonlinear systems. Journal of Mathematical Physics, 5(3):

383–398, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i3/p383_s1.

Leonard:1964:SAA

[560] A. Leonard and T. W. Mullikin. Spectral analysis of the anisotropic neutron transport kernel in slab geometry with applications. Journal of Mathematical Physics, 5(3):399–411, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i3/p399_s1.

Reynolds:1964:SER

[561] J. T. Reynolds, D. S. Onley, and L. C. Biedenharn. Some exact radial integrals for Dirac–Coulomb functions. Journal of Mathematical Physics, 5(3):411–419, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i3/p411_s1.

Olsson:1964:IPD

[562] Per O. M. Olsson. Integration of the partial differential equations for the hypergeometric functions F_{1} and F_{D} of two and more variables. Journal of Mathematical Physics, 5(3):420–430, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i3/p420_s1.

Polkinghorne:1964:CHE

[563] J. C. Polkinghorne. The complete high-energy behavior of ladder diagrams in perturbation theory. Journal of Mathematical Physics, 5(3):431–434, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i3/p431_s1.

Bilhorn:1964:RCR

[564] D. E. Bilhorn, L. L. Foldy, R. M. Thaler, W. Tobocman, and V. A. Madsen. Remarks concerning reciprocity in quantum mechanics. Journal of Mathematical Physics, 5(4):435–441, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i4/p435_s1.

Prugovecki:1964:MSA

[565] Eduard Prugovecki. Multiplicative symmetries in axiomatic quantum field theory. Journal of Mathematical Physics, 5(4):442–455, April 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p442_s1.

Fong:1964:RPD

[566] R. Fong and J. Sucher. Relativistic particle dynamics and the S matrix. Journal of Mathematical Physics, 5(4):456–470, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p456_s1.

Aizu:1964:FOB

[567] Ko Aizu, Gianfausto Dell’Antonio, and Arnold J. F. Siegert. Field operators for bosons with impenetrable cores. II. Equations of motion and general operator formalism. Journal of Mathematical Physics, 5(4):471–489, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i4/p471_s1.

Zeeman:1964:CIL

[568] E. C. Zeeman. Causality implies the Lorentz group. Journal of Mathematical Physics, 5(4):490–493, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i4/p490_s1.

Gabriel:1964:NMR

[569] John Gabriel. New methods for reduction of group representations using an extension of Schur’s lemma. Journal of Mathematical Physics, 5(4): 494–504, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i4/p494_s1.

Mano:1964:RNC

[570] Koichi Mano. Representations for the nonrelativistic Coulomb Green’s function. Journal of Mathematical Physics, 5(4):505–508, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p505_s1.

McCumber:1964:TVSb

[571] D. E. McCumber. Theory of vibrational structure in optical spectra of impurities in solids. II. Multiplets. Journal of Mathematical Physics, 5(4): 508–521, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i4/p508_s1.

Andersen:1964:ECP

[572] H. C. Andersen, I. Oppenheim, Kurt E. Shuler, and George H. Weiss. Exact conditions for the preservation of a canonical distribution in Markovian relaxation processes. Journal of Mathematical Physics, 5(4):522–536, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i4/p522_s1.

Karal:1964:EEO

[573] Frank C. Karal and Joseph B. Keller. Elastic, electromagnetic, and other waves in a random medium. Journal of Mathematical Physics, 5(4):537– 547, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i4/p537_s1.

Keller:1964:TCC

[574] Joseph B. Keller. A theorem on the conductivity of a composite medium. Journal of Mathematical Physics, 5(4):548–549, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p548_s1.

Islam:1964:SEW

[575] Mohammed Azizul Islam. Scattering of electromagnetic waves by a composite cylinder. Journal of Mathematical Physics, 5(4):550–557, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i4/p550_s1.

Silven:1964:AOMa

[576] Saul Silven. Application of operational methods to the analysis of uniform plasmas. Journal of Mathematical Physics, 5(4):557–560, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p557_s1.

See [737].

Curtiss:1964:EIB

[577] C. F. Curtiss. Expansions of integrals of Bessel functions of large order. Journal of Mathematical Physics, 5(4):561–564, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i4/p561_s1.

Lurie:1964:MEO

[578] D. Lurié and A. J. Macfarlane. Matrix elements of the octet operator of SU_{3}. Journal of Mathematical Physics, 5(5):565–575, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p565_s1.

Macfarlane:1964:GSM

[579] A. J. Macfarlane, N. Mukunda, and E. C. G. Sudarshan. Generalized shmushkevich method: Proof of basic results. Journal of Mathematical Physics, 5(5):576–580, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i5/p576_s1.

Streater:1964:IOQ

[580] R. F. Streater. Intensive observables in quantum theory. Journal of Mathematical Physics, 5(5):581–590, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i5/p581_s1.

Hostler:1964:CGF

[581] Levere Hostler. Coulomb Green’s functions and the Furry approximation. Journal of Mathematical Physics, 5(5):591–611, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p591_s1.

Lax:1964:DSS

[582] Peter D. Lax. Development of singularities of solutions of nonlinear hyperbolic partial differential equations. Journal of Mathematical Physics, 5(5): 611–613, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i5/p611_s1.

Balachandran:1964:BCP

[583] A. P. Balachandran. Boundary conditions for a partial-wave amplitude. Journal of Mathematical Physics, 5(5):614–621, May 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i5/p614_s1.

McGuire:1964:SES

[584] J. B. McGuire. Study of exactly soluble one-dimensional N-body problems. Journal of Mathematical Physics, 5(5):622–636, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p622_s1.

Wong:1964:ANF

[585] K. W. Wong. Application of nonlocal field operators to a system of hard-sphere Bose gas. Journal of Mathematical Physics, 5(5):637–642, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i5/p637_s1.

Wertheim:1964:ASP

[586] M. S. Wertheim. Analytic solution of the Percus–Yevick equation. Journal of Mathematical Physics, 5(5):643–651, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i5/p643_s1.

Jones:1964:IIC

[587] Gerald L. Jones. The influence of initial correlations on the approach to equilibrium. Journal of Mathematical Physics, 5(5):651–658, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p651_s1.

Salpeter:1964:MSD

[588] E. E. Salpeter and S. B. Treiman. Multiple scattering in the diffusion approximation. Journal of Mathematical Physics, 5(5):659–668, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p659_s1.

Mendelson:1964:OSN

[589] M. R. Mendelson and G. C. Summerfield. One-speed neutron transport in two adjacent half-spaces. Journal of Mathematical Physics, 5(5):668–674, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/

i5/p668_s1. See addendum [733].

Weiss:1964:SDF

[590] George H. Weiss. Simple derivation of the Faxén solution to the Lamm equation. Journal of Mathematical Physics, 5(5):675–676, May 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p675_s1.

Mehta:1964:PSF

[591] C. L. Mehta. Phase-space formulation of the dynamics of canonical variables. Journal of Mathematical Physics, 5(5):677–686, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p677_s1.

Gross:1964:NIM

[592] Leonard Gross. Norm invariance of mass-zero equations under the conformal group. Journal of Mathematical Physics, 5(5):687–695, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p687_s1.

Lipkin:1964:ENC

[593] Daniel M. Lipkin. Existence of a new conservation law in electromagnetic theory. Journal of Mathematical Physics, 5(5):696–700, May 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i5/p696_s1.

Prosser:1964:SQP

[594] Reese T. Prosser. Segal’s quantization procedure. Journal of Mathematical Physics, 5(6):701–707, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i6/p701_s1.

Prosser:1964:CPE

[595] Reese T. Prosser. Convergent perturbation expansions for certain wave operators. Journal of Mathematical Physics, 5(6):708–713, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p708_s1.

Garczynski:1964:NSC

[596] W. Garczynski. The necessary and sufficient condition in terms of Wightman functions for a field to be a generalized free field. Journal of Mathematical Physics, 5(6):714–719, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i6/p714_s1.

Peres:1964:DWE

[597] Asher Peres. On Dirac’s wave equation in a gravitational field. Journal of Mathematical Physics, 5(6):720–721, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i6/p720_s1.

Ghirardi:1964:PCN

[598] G. C. Ghirardi and A. Rimini. Properties of a class of nonlocal solvable interactions. Journal of Mathematical Physics, 5(6):722–728, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p722_s1.

Dollard:1964:ACC

[599] John D. Dollard. Asymptotic convergence and the Coulomb interaction. Journal of Mathematical Physics, 5(6):729–738, June 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i6/p729_s1.

Secrest:1964:EES

[600] Don Secrest and Lloyd Martin Holm. Eigenfunctions of the electron spin operators. Journal of Mathematical Physics, 5(6):738–742, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p738_s1.

Weinberg:1964:PTS

[601] Steven Weinberg. Perturbation theory for strong repulsive potentials. Journal of Mathematical Physics, 5(6):743–747, June 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i6/p743_s1.

Ron:1964:QKE

[602] Amiram Ron. Quantum kinetic equations for electrons in a periodic system. Journal of Mathematical Physics, 5(6):748–755, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p748_s1.

Leff:1964:SCO

[603] Harvey S. Leff. Systematic characterization of mth-order energy-level spacing distributions. Journal of Mathematical Physics, 5(6):756–762, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i6/p756_s1.

Leff:1964:CES

[604] Harvey S. Leff. Class of ensembles in the statistical theory of energylevel spectra. Journal of Mathematical Physics, 5(6):763–768, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p763_s1.

Soda:1964:SDP

[605] Toshio Soda. Some dynamical properties of an impurity in the hard-sphere Bose gas. Journal of Mathematical Physics, 5(6):769–774, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p769_s1.

McKean:1964:KWD

[606] H. P. McKean. Kramers–Wannier duality for the 2-dimensional Ising model as an instance of Poisson’s summation formula. Journal of Mathematical Physics, 5(6):775–776, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i6/p775_s1.

Roman:1964:RCB

[607] P. Roman. A remarkable connection between Kemmer algebras and unitary groups. Journal of Mathematical Physics, 5(6):777–783, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i6/p777_s1.

Callaway:1964:TSS

[608] Joseph Callaway. Theory of scattering in solids. Journal of Mathematical Physics, 5(6):783–798, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i6/p783_s1.

Pais:1964:SPP

[609] A. Pais and Tai Tsun Wu. Singular potentials and peratization. II. Journal of Mathematical Physics, 5(6):799–804, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i6/p799_s1.

Burke:1964:SSW

[610] J. E. Burke. Scattering of surface waves on an infinitely deep fluid. Journal of Mathematical Physics, 5(6):805–819, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i6/p805_s1.

Chen:1964:DST

[611] Yung Ming Chen. Diffraction by a smooth transparent object. Journal of Mathematical Physics, 5(6):820–832, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i6/p820_s1.

Moses:1964:GJF

[612] H. E. Moses. Generalizations of the Jost functions. Journal of Mathematical Physics, 5(6):833–840, June 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i6/p833_s1.

Lebowitz:1964:CVE

[613] J. L. Lebowitz and O. Penrose. Convergence of virial expansions. Journal of Mathematical Physics, 5(7):841–847, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i7/p841_s1.

Haag:1964:AAQ

[614] Rudolf Haag and Daniel Kastler. An algebraic approach to quantum field theory. Journal of Mathematical Physics, 5(7):848–861, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p848_s1.

Bargmann:1964:NWT

[615] V. Bargmann. Note on Wigner’s theorem on symmetry operations. Journal of Mathematical Physics, 5(7):862–868, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i7/p862_s1.

Korff:1964:ZMR

[616] David Korff. Zero-mass representations of the proper inhomogeneous Lorentz group. Journal of Mathematical Physics, 5(7):869–877, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p869_s1.

McKenna:1964:CRT

[617] James McKenna and John R. Klauder. Continuous-representation theory. IV. Structure of a class of function spaces arising from quantum mechanics. Journal of Mathematical Physics, 5(7):878–896, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p878_s1.

Peaslee:1964:UDS

[618] D. C. Peaslee. Uniqueness of 4- and 8-dimensional spaces. Journal of Mathematical Physics, 5(7):897–899, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i7/p897_s1.

Hagen:1964:LBL

[619] C. R. Hagen. Lower bounds on the Lehmann weights in spin-zero meson theory. Journal of Mathematical Physics, 5(7):900–902, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p900_s1.

Schwartz:1964:LHF

[620] Melvin Schwartz. Lagrangian and Hamiltonian formalisms with supplementary conditions. Journal of Mathematical Physics, 5(7):903–907, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i7/p903_s1.

Swift:1964:SBS

[621] Arthur R. Swift and Benjamin W. Lee. Singular Bethe–Salpeter scattering amplitudes. Journal of Mathematical Physics, 5(7):908–921, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p908_s1.

Chakrabarti:1964:CRK

[622] A. Chakrabarti. On the canonical relativistic kinematics of N-particle systems. Journal of Mathematical Physics, 5(7):922–929, July 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p922_s1.

Lenard:1964:MDG

[623] A. Lenard. Momentum distribution in the ground state of the onedimensional system of impenetrable bosons. Journal of Mathematical Physics, 5(7):930–943, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i7/p930_s1.

Fisher:1964:CFC

[624] Michael E. Fisher. Correlation functions and the critical region of simple fluids. Journal of Mathematical Physics, 5(7):944–962, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p944_s1.

Koppel:1964:SFT

[625] Donald Koppel. On the stability of flow of a thermally stratified fluid under the action of gravity. Journal of Mathematical Physics, 5(7):963–982, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i7/p963_s1.

Gilbert:1964:BIO

[626] R. P. Gilbert. Bergman’s integral operator method in generalized axially symmetric potential theory. Journal of Mathematical Physics, 5(7): 983–997, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i7/p983_s1.

Petschek:1964:FMC

[627] Albert G. Petschek. Fluctuations in multiple capture processes. Journal of Mathematical Physics, 5(7):998–999, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i7/p998_s1.

Soda:1964:EES

[628] Toshio Soda. Errata: “Excitation Spectrum of an Impurity in the Hard-Sphere Bose Gas”. Journal of Mathematical Physics, 5(7):1000, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i7/p1000_s1.

See [530].

Calais:1964:ICM

[629] Jean-Louis Calais and Klaus Appel. Inversion of cyclic matrices. Journal of Mathematical Physics, 5(8):1001–1008, August 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1001_s1.

Stephenson:1964:IMS

[630] John Stephenson. Ising-model spin correlations on the triangular lattice. Journal of Mathematical Physics, 5(8):1009–1024, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1009_s1.

Grossmann:1964:NHS

[631] Alexander Grossmann. Nested Hilbert spaces in quantum mechanics. I. Journal of Mathematical Physics, 5(8):1025–1037, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1025_s1.

Fairbairn:1964:FDS

[632] W. M. Fairbairn, T. Fulton, and W. H. Klink. Finite and disconnected subgroups of SU_{3} and their application to the elementary-particle spectrum. Journal of Mathematical Physics, 5(8):1038–1051, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1038_s1.

Shpiz:1964:PCC

[633] J. M. Shpiz. On Pais’s charge correlation coefficients. Journal of Mathematical Physics, 5(8):1052, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i8/p1052_s1.

Bolsterli:1964:RSA

[634] M. Bolsterli. Renormalization of singlet amplitude in intermediate-vector-boson theory of weak interactions. Journal of Mathematical Physics, 5 (8):1053–1054, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i8/p1053_s1.

Delves:1964:ULBa

[635] L. M. Delves. Upper and lower bounds on generalized Fourier coefficients. Journal of Mathematical Physics, 5(8):1055–1060, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1055_s1.

Delves:1964:ULBb

[636] L. M. Delves. Upper and lower bounds on the matrix elements of an arbitrary bounded operator. Journal of Mathematical Physics, 5(8):1061– 1064, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i8/p1061_s1.

Cox:1964:MCB

[637] Joseph R. Cox. Many-channel Bargmann potentials. Journal of Mathematical Physics, 5(8):1065–1069, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i8/p1065_s1.

Belinfante:1964:ESS

[638] Johan G. Belinfante. Existence of scattering solutions for the Schrödinger equation. Journal of Mathematical Physics, 5(8):1070–1074, August 1964.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1070_s1.

Jackson:1964:TTB

[639] Julius L. Jackson and Sam R. Coriell. On trapped trajectories in Brownian motion. Journal of Mathematical Physics, 5(8):1075–1078, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1075_s1.

Ezawa:1964:RCV

[640] Hiroshi Ezawa. The representation of canonical variables as the limit of infinite space volume: The case of the BCS model. Journal of Mathematical Physics, 5(8):1078–1090, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i8/p1078_s1.

Katsura:1964:LHM

[641] Shigetoshi Katsura and Sakari Inawashiro. Linear Heisenberg model of ferro- and antiferromagnetism. Journal of Mathematical Physics, 5(8): 1091–1106, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i8/p1091_s1.

Rushbrooke:1964:TRD

[642] G. S. Rushbrooke. On the theory of randomly dilute Ising and Heisenberg ferromagnetics. Journal of Mathematical Physics, 5(8):1106–1116, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i8/p1106_s1.

Sykes:1964:ECP

[643] M. F. Sykes and J. W. Essam. Exact critical percolation probabilities for site and bond problems in two dimensions. Journal of Mathematical Physics, 5(8):1117–1127, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i8/p1117_s1.

Kesten:1964:NSA

[644] Harry Kesten. On the number of self-avoiding walks. II. Journal of Mathematical Physics, 5(8):1128–1137, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i8/p1128_s1.

Sherman:1964:PPC

[645] S. Sherman. Product property and cluster property equivalence. Journal of Mathematical Physics, 5(8):1137–1139, August 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1137_s1.

Balescu:1964:KEI

[646] R. Balescu and A. Kuszell. Kinetic equation for an inhomogeneous plasma far from equilibrium. Journal of Mathematical Physics, 5(8):1140– 1149, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i8/p1140_s1.

Glasser:1964:NES

[647] M. L. Glasser. Note on the evaluation of some Fermi integrals. Journal of Mathematical Physics, 5(8):1150–1152, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i8/p1150_s1.

Newman:1964:ESG

[648] David S. Newman. Equation of state for a gas with a weak, long-range positive potential. Journal of Mathematical Physics, 5(8):1153–1157, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i8/p1153_s1.

Fischer:1964:MSI

[649] Gaston Fischer. Microwave surface impedance of a periodic medium. Journal of Mathematical Physics, 5(8):1158–1167, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1158_s1.

Odeh:1964:IEN

[650] Farouk M. Odeh. On the integrodifferential equations of the nonlocal theory of superconductivity. Journal of Mathematical Physics, 5(8):1168– 1182, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i8/p1168_s1.

Schmidt:1964:AEC

[651] Paul W. Schmidt. Asymptotic expansion of certain integrals containing the Bessel function J_{0}(x). Journal of Mathematical Physics, 5(8):1183– 1184, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i8/p1183_s1.

Spector:1964:ESS

[652] Richard M. Spector. Exact solution of the Schrödinger equation for inverse fourth-power potential. Journal of Mathematical Physics, 5(9):1185– 1189, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i9/p1185_s1.

Nelson:1964:INP

[653] Edward Nelson. Interaction of nonrelativistic particles with a quantized scalar field. Journal of Mathematical Physics, 5(9):1190–1197, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i9/p1190_s1.

Lumley:1964:MVR

[654] J. L. Lumley. On multiple-valued random functions. Journal of Mathematical Physics, 5(9):1198–1203, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i9/p1198_s1.

Klauder:1964:LRS

[655] John R. Klauder. Linear representation of spinor fields by antisymmetric tensors. Journal of Mathematical Physics, 5(9):1204–1214, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i9/p1204_s1.

Griffiths:1964:PFE

[656] Robert B. Griffiths. A proof that the free energy of a spin system is extensive. Journal of Mathematical Physics, 5(9):1215–1222, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1215_s1.

Manning:1964:TDS

[657] Irwin Manning. A theorem on the determinantal solution of the Fredholm equation. Journal of Mathematical Physics, 5(9):1223–1225, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i9/p1223_s1.

Antoine:1964:CIRa

[658] J.-P. Antoine and D. Speiser. Characters of irreducible representations of the simple groups. I. General theory. Journal of Mathematical Physics, 5(9):1226–1234, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i9/p1226_s1.

Hostler:1964:NCG

[659] Levere Hostler. Nonrelativistic Coulomb Green’s function in momentum space. Journal of Mathematical Physics, 5(9):1235–1240, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1235_s1.

Willis:1964:MIR

[660] Charles R. Willis. A model of interacting radiation and matter. Journal of Mathematical Physics, 5(9):1241–1252, September 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1241_s1.

Derrick:1964:CNW

[661] G. H. Derrick. Comments on nonlinear wave equations as models for elementary particles. Journal of Mathematical Physics, 5(9):1252–1254, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i9/p1252_s1.

Chow:1964:STP

[662] Yutze Chow. On some topological properties of Feynman graphs and their application to formulas related to the Feynman amplitudes. Journal of Mathematical Physics, 5(9):1255–1260, September 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1255_s1.

Kalikstein:1964:SEW

[663] Kalman Kalikstein and Larry Spruch. Scattering of electromagnetic waves by a ferrite in a waveguide. Journal of Mathematical Physics, 5(9):1261– 1272, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i9/p1261_s1.

Su:1964:KTW

[664] C. H. Su. Kinetic theory of a weakly coupled gas. Journal of Mathematical Physics, 5(9):1273–1290, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i9/p1273_s1.

Kraut:1964:UWS

[665] Edgar A. Kraut. The uniqueness of weak solutions of the one-dimensional scalar analog to the Navier–Stokes equation. Journal of Mathematical Physics, 5(9):1290–1292, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i9/p1290_s1.

Bailey:1964:EQR

[666] Paul B. Bailey. Exact quantization rules for the one-dimensional Schrödinger equation with turning points. Journal of Mathematical Physics, 5(9):1293–1297, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v5/i9/p1293_s1.

Tiersten:1964:CME

[667] H. F. Tiersten. Coupled magnetomechanical equations for magnetically saturated insulators. Journal of Mathematical Physics, 5(9):1298–1318, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i9/p1298_s1.

Yao:1964:QEA

[668] York-Peng Yao. Quantization of electrodynamics in the axial gauge. Journal of Mathematical Physics, 5(9):1319–1321, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1319_s1.

Yao:1964:SSE

[669] York-Peng Yao. Spin and statistics with an electromagnetic field. Journal of Mathematical Physics, 5(9):1322–1326, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i9/p1322_s1.

Seto:1964:WEG

[670] Y. J. Seto and Arwin A. Dougal. The wave equation and the Green’s dyadic for bounded magnetoplasmas. Journal of Mathematical Physics, 5(9):1326–1334, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i9/p1326_s1.

Hagen:1964:TTG

[671] C. R. Hagen and A. J. Macfarlane. Triality type and its generalization in unitary symmetry theories. Journal of Mathematical Physics, 5(9): 1335–1339, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i9/p1335_s1.

Kenschaft:1964:SSI

[672] R. P. Kenschaft and R. D. Amado. Solution of a singular integral equation from scattering theory. Journal of Mathematical Physics, 5(9):1340– 1342, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i9/p1340_s1.

Coleman:1964:CGS

[673] Sidney Coleman. The Clebsch–Gordan series for SU(3). Journal of Mathematical Physics, 5(9):1343–1344, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v5/i9/p1343_s1. See erratum [738].

Jordan:1964:LIM

[674] Thomas F. Jordan. Lorentz invariant multichannel scattering formalism. Journal of Mathematical Physics, 5(10):1345–1360, October 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1345_
s1.

Schroer:1964:CNF

[675] B. Schroer. The concept of nonlocalizable fields and its connection with nonrenormalizable field theories. Journal of Mathematical Physics, 5(10): 1361–1367, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i10/p1361_s1.

Brulin:1964:GSP

[676] Olof Brulin and Stig Hjalmars. The graviton as a spin-2 particle. Journal of Mathematical Physics, 5(10):1368–1370, October 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1368_s1.

Weinberg:1964:HID

[677] Steven Weinberg. The hose instability dispersion relation. Journal of Mathematical Physics, 5(10):1371–1386, October 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1371_s1.

Bardotti:1964:MCC

[678] Graziella Bardotti, Bruno Bertotti, and Laura Gianolio. Magnetic configuration of a cylinder with infinite conductivity. Journal of Mathematical Physics, 5(10):1387–1390, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i10/p1387_s1.

Andrews:1964:CAM

[679] M. Andrews and J. Gunson. Complex angular momenta and many-particle states. I. Properties of local representations of the rotation group. Journal of Mathematical Physics, 5(10):1391–1400, October 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1391_s1.

Morita:1964:SMQ

[680] Tohru Morita. Statistical mechanics of quenched solid solutions with application to magnetically dilute alloys. Journal of Mathematical Physics, 5(10):1401–1405, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i10/p1401_s1.

Jin:1964:RPB

[681] Y. S. Jin and A. Martin. Remarks on the polynomial boundedness in the Mandelstam representation. Journal of Mathematical Physics, 5(10): 1406–1412, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i10/p1406_s1.

Fikioris:1964:MSW

[682] J. G. Fikioris and P. C. Waterman. Multiple scattering of waves. II. “hole corrections” in the scalar case. Journal of Mathematical Physics, 5(10): 1413–1420, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/

1/jmapaq/v5/i10/p1413_s1.

Ashour:1964:IEA

[683] Attia A. Ashour. An integral equation for the associated Legendre function of the first kind. Journal of Mathematical Physics, 5(10):1421–1423, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i10/p1421_s1.

Silven:1964:AOMb

[684] Saul Silven. Application of operational methods to the analysis of the motion of rigid bodies. Journal of Mathematical Physics, 5(10):1424– 1426, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i10/p1424_s1.

Bander:1964:ISD

[685] Myron Bander. An iterative solution of the N/D equation. Journal of Mathematical Physics, 5(10):1427–1430, October 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1427_s1.

Penney:1964:DIR

[686] R. Penney. Duality invariance and Riemannian geometry. Journal of Mathematical Physics, 5(10):1431–1437, October 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i10/p1431_s1.

Lomont:1964:RILb

[687] J. S. Lomont and H. E. Moses. Representations of the inhomogeneous Lorentz group in terms of an angular momentum basis: Derivation for the cases of nonzero mass and zero mass, discrete spin. Journal of Mathematical Physics, 5(10):1438–1457, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i10/p1438_s1. See [735].

Nakanishi:1964:FPP

[688] Noboru Nakanishi. Fundamental properties of perturbation-theoretical integral representations. III. Journal of Mathematical Physics, 5(10):1458– 1473, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i10/p1458_s1.

Granzow:1964:DTO

[689] Kenneth D. Granzow. N-dimensional total orbital angular-momentum operator. II. Explicit representations. Journal of Mathematical Physics, 5(10):1474–1477, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i10/p1474_s1.

Falk:1964:IRN

[690] Harold Falk. Inequalities relating the nearest-neighbor spin correlation and the magnetization for the Heisenberg Hamiltonian. Journal of Mathematical Physics, 5(10):1478–1480, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i10/p1478_s1.

Muzinich:1964:CMH

[691] Ivan J. Muzinich. Crossing matrices for helicity amplitudes, application to crossed channel partial-wave analysis, and Reggeization. Journal of Mathematical Physics, 5(11):1481–1490, November 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1481_s1.

Polkinghorne:1964:ABF

[692] J. C. Polkinghorne. Asymptotic behavior of Feynman integrals with spin. Journal of Mathematical Physics, 5(11):1491–1498, November 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1491_
s1.

Odeh:1964:PDE

[693] Farouk Odeh and Joseph B. Keller. Partial differential equations with periodic coefficients and Bloch waves in crystals. Journal of Mathematical Physics, 5(11):1499–1504, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i11/p1499_s1.

Rosen:1964:MWA

[694] Mervine Rosen and Donald R. Yennie. A modified WKB approximation for phase shifts. Journal of Mathematical Physics, 5(11):1505–1515, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i11/p1505_s1.

Leibowitz:1964:PMR

[695] Martin A. Leibowitz. Particle moderation: Random functional approach. Journal of Mathematical Physics, 5(11):1516–1526, November 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1516_
s1.

Bowden:1964:SIV

[696] Robert L. Bowden and Clayton D. Williams. Solution of the initial-value transport problem for monoenergetic neutrons in slab geometry. Journal of Mathematical Physics, 5(11):1527–1540, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1527_s1.

Chen:1964:SWO

[697] Yung Ming Chen. On scattering of waves by objects imbedded in random media: Stochastic linear partial differential equations and scattering of waves by conducting sphere imbedded in random media. Journal of Mathematical Physics, 5(11):1541–1546, November 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1541_s1.

Redmond:1964:SRC

[698] Peter J. Redmond. Some remarks concerning a pathological matrix of interest in the inverse-scattering problem. Journal of Mathematical Physics, 5(11):1547–1554, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i11/p1547_s1.

Montet:1964:ERP

[699] George L. Montet. The effective resistance of passive networks. Journal of Mathematical Physics, 5(11):1555–1559, November 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1555_s1.

Antoine:1964:CIRb

[700] J.-P. Antoine and D. Speiser. Characters of irreducible representations of the simple groups. II. Application to classical groups. Journal of Mathematical Physics, 5(11):1560–1572, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i11/p1560_s1.

Hashitsume:1964:KAT

[701] N. Hashitsume and S. Fujita. Kinetic approach to thermal transport phenomena. I. Classical theory. Journal of Mathematical Physics, 5(11):1572– 1579, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i11/p1572_s1.

Snider:1964:PVM

[702] R. F. Snider. Perturbation variation methods for a quantum Boltzmann equation. Journal of Mathematical Physics, 5(11):1580–1587, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i11/p1580_s1.

Davidon:1964:ORQ

[703] W. C. Davidon and H. Ekstein. Observables in relativistic quantum mechanics. Journal of Mathematical Physics, 5(11):1588–1594, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i11/p1588_s1.

Gluckstern:1964:RCS

[704] R. L. Gluckstern and Shin-R Lin. Relativistic Coulomb scattering of electrons. Journal of Mathematical Physics, 5(11):1594–1602, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1594_
s1.

Lubkin:1964:PRB

[705] Elihu Lubkin. Possible relationship between electric charge and dual charge. Journal of Mathematical Physics, 5(11):1603–1606, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i11/p1603_s1.

Schwinger:1964:CGF

[706] Julian Schwinger. Coulomb Green’s function. Journal of Mathematical Physics, 5(11):1606–1608, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i11/p1606_s1.

Roos:1964:CUM

[707] Matts Roos. On the construction of a unitary matrix with elements of given moduli. Journal of Mathematical Physics, 5(11):1609–1611, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/

i11/p1609_s1. See erratum [736].

Smith:1964:GAT

[708] G. F. Smith. On the generation of anisotropic tensors. Journal of Mathematical Physics, 5(11):1612–1614, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i11/p1612_s1.

Levine:1964:LFP

[709] Arnold D. Levine. Lagrangian formulation of the phonon field equations. Journal of Mathematical Physics, 5(11):1615–1618, November 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1615_
s1.

Mathur:1964:MSE

[710] N. C. Mathur and K. C. Yeh. Multiple scattering of electromagnetic waves by random scatterers of finite size. Journal of Mathematical Physics, 5 (11):1619–1628, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i11/p1619_s1.

Estabrook:1964:DAS

[711] Frank B. Estabrook and Hugo D. Wahlquist. Dyadic analysis of space-time congruences. Journal of Mathematical Physics, 5(11):1629–1644, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i11/p1629_s1.

Fradkin:1964:SWD

[712] D. M. Fradkin, C. L. Hammer, and T. A. Weber. Scattering wavefunctions for a Dirac particle in a central potential. Journal of Mathematical Physics, 5(11):1645–1651, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i11/p1645_s1.

Barut:1964:CLG

[713] A. O. Barut. Complex Lorentz group with a real metric: Group structure. Journal of Mathematical Physics, 5(11):1652–1656, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i11/p1652_
s1.

Penney:1964:ADB

[714] R. Penney. Algebra of Dirac bilinears. Journal of Mathematical Physics, 5(11):1657–1658, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i11/p1657_s1.

Morgan:1964:TCN

[715] Thomas A. Morgan. Two classes of new conservation laws for the electromagnetic field and for other massless fields. Journal of Mathematical Physics, 5(11):1659–1660, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i11/p1659_s1.

Bronk:1964:TRE

[716] Burt V. Bronk. Theorem relating the eigenvalue density for random matrices to the zeros of the classical polynomials. Journal of Mathematical Physics, 5(12):1661–1663, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i12/p1661_s1.

Corben:1964:IMW

[717] H. C. Corben. If microns were fermis. Journal of Mathematical Physics, 5(12):1664–1668, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i12/p1664_s1.

Lee:1964:ERP

[718] K. S. H. Lee and C. H. Papas. Electromagnetic radiation in the presence of moving simple media. Journal of Mathematical Physics, 5(12):1668– 1672, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i12/p1668_s1.

Morrison:1964:HFMa

[719] J. A. Morrison and H. L. Frisch. High-field magnetoresistance of inhomogeneous semiconductors and plasmas. II. Two-dimensional inhomogeneity distributions. Journal of Mathematical Physics, 5(12):1672–1681, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i12/p1672_s1.

Morrison:1964:HFMb

[720] J. A. Morrison and H. L. Frisch. High-field magnetoresistance of inhomogeneous semiconductors and plasmas. III. Two-dimensional inhomogeneity distributions. Journal of Mathematical Physics, 5(12):1681–1687, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i12/p1681_s1.

Isenberg:1964:NME

[721] C. Isenberg. Numerical method for the exact expansion of generating functions. Journal of Mathematical Physics, 5(12):1687–1691, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1687_
s1.

Kostin:1964:GVM

[722] Morton D. Kostin and Harvey Brooks. Generalization of the variational method of kahan, rideau, and roussopoulos and its application to neutron transport theory. Journal of Mathematical Physics, 5(12):1691–1700, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i12/p1691_s1.

Wu:1964:UOP

[723] Ching-Sheng Wu. Useful operator in plasma kinetic theory. Journal of Mathematical Physics, 5(12):1701–1712, December 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1701_s1.

Tarski:1964:RFT

[724] Jan Tarski. Representation of fields in a two-dimensional model theory. Journal of Mathematical Physics, 5(12):1713–1722, December 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1713_
s1.

Baird:1964:RSLa

[725] G. E. Baird and L. C. Biedenharn. On the representations of the semisimple Lie groups. III. The explicit conjugation operation for SU_{n}. Journal of Mathematical Physics, 5(12):1723–1730, December 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1723_s1.

Baird:1964:RSLb

[726] G. E. Baird and L. C. Biedenharn. On the representations of the semisimple Lie groups. IV. A canonical classification for tensor operators in SU_{3}. Journal of Mathematical Physics, 5(12):1730–1747, December 1964. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i12/p1730_
s1.

Chakrabarti:1964:ALT

[727] A. Chakrabarti. Applications of the Lorentz transformation properties of canonical spin tensors. Journal of Mathematical Physics, 5(12):1747– 1755, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i12/p1747_s1.

Garrod:1964:RPV

[728] Claude Garrod and Jerome K. Percus. Reduction of the N-particle variational problem. Journal of Mathematical Physics, 5(12):1756–1776, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i12/p1756_s1.

Mitzner:1964:ESI

[729] K. M. Mitzner. Effect of small irregularities on electromagnetic scattering from an interface of arbitrary shape. Journal of Mathematical Physics, 5(12):1776–1786, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i12/p1776_s1.

Borsch-Supan:1964:CTM

[730] Wolfgang Börsch-Supan. Comparison of two methods for lower bounds to eigenvalues. Journal of Mathematical Physics, 5(12):1787–1788, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/
i12/p1787_s1.

Montgomery:1964:GPE

[731] David Montgomery. Generalized perturbation expansion for the Klein– Gordon equation with a small nonlinearity. Journal of Mathematical Physics, 5(12):1788–1795, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i12/p1788_s1.

Gruber:1964:TCI

[732] B. Gruber and L. O’Raifeartaigh. S theorem and construction of the invariants of the semisimple compact Lie algebras. Journal of Mathematical Physics, 5(12):1796–1804, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v5/i12/p1796_s1.

Mendelson:1964:AOS

[733] M. R. Mendelson and G. C. Summerfield. Addendum: “One-Speed Neutron Transport in Two Adjacent Half-Spaces”. Journal of Mathematical Physics, 5(12):1804, December 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i12/p1804_s1. See [589].

Lomont:1965:ERIa

[734] J. S. Lomont and H. E. Moses. Errata: “The Representations of the Inhomogeneous Lorentz Group in Terms of an Angular Momentum Basis”. Journal of Mathematical Physics, 6(4):670, April 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i4/p670_s1.

Lomont:1965:ERIb

[735] J. S. Lomont and H. E. Moses. Errata: “Representations of the Inhomogeneous Lorentz Group in Terms of an Angular Momentum Basis: Derivation for the Cases of Nonzero Mass and Zero Mass, Discrete Spin”. Journal of Mathematical Physics, 6(4):670, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i4/p670_s2. See [687].

Roos:1965:ECU

[736] Matts Roos. Erratum: “On the Construction of a Unitary Matrix with Elements of Given Moduli”. Journal of Mathematical Physics, 6(8): 1354, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1354_s1. See [707].

Silven:1965:EAO

[737] Saul Silven. Errata: “Application of Operational Methods to The Analysis of Uniform Plasmas”. Journal of Mathematical Physics, 6(8):1354, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/
i8/p1354_s2. See [576].

Coleman:1965:ECG

[738] Sidney Coleman. Erratum: “The Clebsch–Gordan Series for SU(3)”. Journal of Mathematical Physics, 6(12):2014, December 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i12/p2014_s1. See [673].

Glasser:1966:ENE

[739] M. L. Glasser. Erratum: “Note on the Evaluation of Some Fermi Integrals”. Journal of Mathematical Physics, 7(7):1340, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1340_s1. See

[647].

Yeh:1988:EDW

[740] C. Yeh and C. S. Kim. Erratum: “The diffraction of waves by a penetrable ribbon”, [J. Math. Phys. 4 (1963), 65–71; MR 26 #1023] by Yeh. Journal of Mathematical Physics, 29(3):721–722, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i3/p721_s1. See [333].